### v1

# MMAO-Bench: MultiModal All in One Benchmark Reveals Compositional Law between Uni-modal and Omni-modal in OmniModels

Abstract
Multimodal Large Languages models have been progressing from uni-modal understanding toward unifying visual, audio and language modalities, collectively termed omni models. However, the correlation between uni-modal and omni-modal remains unclear, which requires comprehensive evaluation to drive omni model’s intelligence evolution. In this work, we propose a novel, high quality and diversity omni model benchmark, MultiModal All in One Benchmark (MMAO-Bench), which effectively assesses both uni-modal and omni-modal understanding capabilities. The benchmark consists of 1880 human curated samples, across 44 task types, and a innovative multi-step open-ended question type that better assess complex reasoning tasks. Experimental result shows the compositional law between cross-modal and uni-modal performance and the omni-modal capability manifests as a bottleneck effect on weak models, while exhibiting synergistic promotion on strong models.
MMAO-Bench: MultiModal All in One Benchmark Reveals Compositional Law between Uni-modal and Omni-modal in OmniModels
Chen Chenthanks: Equal contribution.   ZeYang Hu11footnotemark: 1   Fengjiao Chenthanks: Project Lead. Liya Ma   Jiaxing Liu   Xiaoyu Li22footnotemark: 2   Xuezhi Cao Meituan, Beijing, China. {chenchen165, huzeyang, chenfengjiao02, maliya06}@meituan.com {liujiaxing10, lixiaoyu28, caoxuezhi}@meituan.com
1 Introduction
Multimodal artificial intelligence has undergone extensive researches in visual language model and audio language model, with current advancements progressing toward unifying visual, audio and language modalities, collectively termed omni models. The evaluation paradigm for these models has consequently expanded from assessing uni-modal understanding capabilities (i.e. visual understanding, audio understanding) to the next-level of intelligence (i.e. omni-modal understanding).
Existing omni-model evaluation benchmarks remain relatively scarce and focus on different aspects. For example, some prioritize image comprehension(Li et al., 2024b), others emphasize video understanding(Hong et al., 2025), while a subset concentrates on speech interaction(Gong et al., 2024). Notably, existing datasets are exclusively English-centric, lacking evaluation benchmarks for Chinese linguistic contexts.
The ideal omni-model should simultaneously preserve visual understanding capabilities (e.g., MMBench(Liu et al., 2024a)/MathVista(Lu et al., 2024)/MVBench(Li et al., 2024a)), speech interaction proficiency (e.g., MMAU(Sakshi et al., 2025)), and cross-modal integration capacity (e.g., OmniBench(Li et al., 2024b)/WorldSense(Hong et al., 2025)). Current evaluation paradigms employ disjointed benchmark suites for separate capability testing, creating resource-intensive evaluation processes and disconnected modality assessments. Beyond uni-modal, omni-modal capability introduces advanced challenges across image, video and audio modality. However, 77% questions from WorldSense are solvable without video or audio, and 25% questions from OmniBench contain erroneous answers. These designs limit the evaluation and analysis of omni models’ abilities.
Addressing theses limitations, we propose a novel, high quality and diversity omni model benchmark, MultiModal All in One Benchmark (MMAO-Bench), that enables comprehensive model assessment and pushing omni model to the next-level of intelligence.
Main Contributions:
1. Propose the first MultiModal All in One Benchmark evaluation framework specifically designed for omni models, which efficiently assesses both uni-modal and omni-modal understanding capabilities. This framework verifies the compositional law between omni-modal and uni-modal capability. For weaker models, this manifests itself as a "short-board effect" (where performance is constrained by the weakest modality), while stronger models demonstrate emergent ability across modalities.
2. Establish a systematic and unified dataset framework, featuring complete manual annotation to ensure high quality and diversity. This approach effectively prevents data contamination while better aligning with practical application scenarios. As a result, the MMAO-Bench comprises 44 different types of tasks and is 100% passing the quality check, which significantly surpassing existing datasets.
3. Beyond conventional multiple-choice formats, the evaluation incorporates innovative Multi-Step Open-Ended Question tasks to show a more realistic and discriminative evaluation result on complex reasoning.
2 Related Work
2.1 Uni-Modal Benchmarks
Based on large language models, vision language models (VLM) (Bai et al., 2025; Xiaomi, 2025; Zeng et al., 2025) and audio language models (ALM) (Ding et al., 2025; Wu et al., 2025) introduce the general intelligence to vision modality and audio modality respectively. Various uni-modal benchmarks conduct comprehensive evaluations on VLMs (Liu et al., 2024a; Lu et al., 2024; Wang et al., 2024a, b; Liu et al., 2024c; Mathew et al., 2021; Ouyang et al., 2024; Li et al., 2024a; Wu et al., 2024; Liu et al., 2024b; xAI, 2023; Xiao et al., 2021; Huang et al., 2025; Hu et al., 2025; Fu et al., 2024) and ALMs (Ardila et al., 2019; Wang et al., 2021; Yang et al., 2024; Ao et al., 2024). For image modality, MMBench(Liu et al., 2024a) proposed a systematically designed benchmark to evaluate general image understanding on 20 different tasks. Focused on mathematic, MathVision(Wang et al., 2024a) collected questions from 19 mathematic competitions to evaluate VLMs complex reasoning ability. In addition to above, OCRBench(Liu et al., 2024c) supplied the evaluation on text recognition and document understanding. For video modality, MVBench(Li et al., 2024a) aggregated 11 public video benchmarks and incorporated data enhancement process to cover 20 dynamic video understanding tasks. To complement the long video understanding area, LongVideoBench(Wu et al., 2024) introduces hourly video materials to evaluate the information retrieval ability from long context. For audio modality, MMAU(Sakshi et al., 2025) provides general audio understanding assessment across speech, sounds and music domains, featuring diverse audio samples. There’s massive uni-modal benchmarks covering diverse model abilities on vision modality and audio modality separately.
2.2 Omni-Modal Benchmarks
Omni models have arisen in recent years(Comanici et al., 2025; Xu et al., 2025b; AI et al., 2025; Li et al., 2025), as the pioneer, Gemini(Comanici et al., 2025) shows a strong ability in understanding both vision and audio, while qwen-3-omni(Xu et al., 2025b) provides leading performance in open-sourced models. However, there’s less omni-modal benchmarks that can evaluate the modality combination across image, video and audio. OmniBench(Li et al., 2024b) inserted audio as a context into the image understanding task and made up an omni-modal benchmark, while the data quality needs further improvement. WorldSence(Hong et al., 2025) emphasized audio-visual data in real world scenarios with high data quality, while most audio-visual questions can be solved by audio or video alone, which can’t assess the cross-modality ability. Other datasets focus on audio (Gong et al., 2024) or video (Zhou et al., 2025) and cover limited task types.
Due to the limited quality and coverage of existing benchmarks, we introduce a new benchmark MMAO-Bench which comprise 1250 human crafted questions for omni-modal and 630 enhanced samples for uni-modal, across 44 task types with 98% cross-modality solvable. The materials are collected from in-house dataset with manual annotation which prevents data contamination while better aligning with practical application scenarios. Beyond conventional multiple-choice formats, the evaluation incorporates innovative multi-step open-ended question tasks to show a more realistic and discriminative evaluation result on complex reasoning. Besides the self-constructed omni-modal dataset, we construct a uni-modal dataset by aggregating public uni-model datasets systematically and design a clustering-guided sampling method to achieves both evaluation efficiency and consistency. In this way, our benchmark can have a thoroughly evaluation that requires omni models to keep uni-modal abilities and obtain omni-modal abilities at the same time.
Dataset Omni-modal Uni-modal Acc. Solvable Source #Tasks #QA Pairs QA Type Language
MMBench I - - 80%private 20 3217 MC EN/ZH
MMAU A - - 15%private 27 10000 MC EN
MVBench V - - public 20 4000 MC EN
OmniBench I+A 75% 90% public 8 1142 MC EN
AV-Odyssey I+V+A 91% 99% public 26 4555 MC EN
WorldSense V+A 99% 23% public 26 3172 MC EN
Daily-Omni V+A - - public 6 1197 MC EN
MMAO-Bench I+V+A I/V/A 100% 98% 90%private 44 1880 MC/MO EN/ZH
Table 1: Comparison of MultiModal Benchmarks. Where I,A,V and T indicate image, audio, video and text modality. The accuracy of question answer pairs and the percent of questions needing omni-modal to solve is shown in Acc. and Solvable. Source indicates where the materials are collected from, public datasets or private datasets that can prevent data contamination. For QA type, MC, SO and MO indicates multi-choise QA, single-open question QA and multi-open question QA. EN and ZH indicates english and chinese.
3 Method
In this section, we first introduce the comprehensive design of model ability system in Section.3.1. The the high-quality data construction pipeline are introduced in Section.3.2. The multi-step open-ended questions are introduced in Section.3.4. Finally, to improve the evaluation efficiency, we introduce a general evaluation dataset compression method in Section.3.3.2.
3.1 Model Ability System
As shown in Figure.1, the model’s capabilities are systematically categorized into two primary dimensions: Perception and Reasoning.
Perception Layer Structured through six recognition types: Object Perception, Attribute Perception, Scenario Perception, Spatial Perception, Cross- Modal Conversion, Semantic Understanding. In addition, we incorporate Cross-Modal Alignment evaluation to assess information synchronization across modalities.
Reasoning Layer extends conventional reasoning categories (General, STEM, Code) with: Spatial Reasoning, Temporal Reasoning, and Complex Reasoning (multi-conditional, multi-step problem solving).
This ability system combines uni-modal and cross-modal abilities which enables cross-modal measurement that is particularly critical for advanced omni-models. Based on this system, we create a diversity dataset with 44 task types distributed as shown in Figure.2.
Figure 1: Illustration of the omni-modal ability system proposed in MMAO-Bench.
Figure 2: Data distribution in model abilities and corresponding task types.
3.2 Omni-modal Dataset Construction
We have established a standardized and efficient data construction process (Figure.3) that produces high-quality and high-diversity dataset.
Figure 3: Data Construction Pipeline
Preparation: (1) Task Design. Deconstruct cross-modal capability systems and task types. Details can be seen in Figure.2. (2) Material Collection. Our materials have the following three characteristics: a. Diverse Sources. The majority of our materials are real-world photos and videos collected through crowdsourcing, with another portion sourced from copyright-free website. Additionally, a small fraction comes from high-quality public datasets such as MMVU, LongVideoBench, and VideoVista. b. Rich and Diverse Topics. Our materials cover a broad spectrum of subjects, including society, culture, art, life, literature, science, and so on. c. Live-Recorded Audio. Apart from background sounds and music, all dialogue is recorded by human speakers. With over 20 participants in the recording process, the audio features are rich and closely reflect the diverse vocal characteristics of the real world. (3) Material filtering. Eliminate meaningless, illogical, and low-quality materials, and categorize the remaining materials by theme to create a material library. Additionally, label the materials with more detailed information such as subject, event, scene, and style to facilitate subsequent annotators to quickly find matching materials.
QA Annotation: (1) Annotators. It mainly includes two parts: human experts and high-quality crowd-sourced users. Human experts have extensive experience in cross-modal data construction and annotation, a deeper understanding of model capabilities, and thus ensure higher professionalism and specificity in the data they construct. Most crowd-sourced users are college students with rich experience in multi-modal model interactions and diverse professional backgrounds, providing data with better authenticity and diversity. (2) Annotation process. First, annotators clarify the required image/video features based on task type definitions and filter appropriate materials from existing libraries using tags. Second, following data construction requirements, they then design prompts and corresponding answers. Third, to enhance data authenticity, all dialogue audio is recorded manually. Through this workflow, we ultimately generate complete QA pairs encompassing three modalities: visual, auditory, and textual.
Quality Inspection: We have established a multi-stage, cyclically validated quality assurance system composed of automated tools and manual review. Each question undergoes at least three rounds of independent quality inspection to maximize data quality. First, a preliminary model check is conducted to filter out cases with ambiguous questions, non-unique answers, or those that do not conform to the task type. Secondly, through modality ablation experiments, we remove one modality of information from the QA pair to see if the model can answer based solely on the remaining information. If the question becomes unanswerable or ambiguous after removing any one modality, it proves the three-modality solvability of the data. Finally, manual quality inspection and revision are performed.
3.3 Uni-modal Dataset Improvement
3.3.1 Quality Improvement
Existing public uni-modal datasets are bothered by data leakage issue. To verify the influence, we adopt privatization improvement on the widely used public dataset MMBench. As shown in Figure4, experiments show that the performance of models have better distinguishability after dataset improvement, reflecting the true capability differences between models. Therefore, for uni-modal data, we also follow the aforementioned construction process for self-construction datasets.
Figure 4: Comparison result of privatization improvement. After improvement, the performances among models are more progressively and distinguishable.
In addition to self-constructed data, we also selected some multimodal data from public datasets to supplement in terms of capability items and data types. (Data mainly comes from AV-Odyssey and WorldSense, accounting for 11% of the total). The specific selection logic is as follows:
1. Comprehensiveness: In terms of capabilities, focus on supplementing the perception part with a relatively low self-construction proportion, while also adding some reasoning questions; in terms of data types, prioritize selecting video + audio modality combinations with a lower self-construction proportion for supplementation, followed by image + audio.
2. Diversity: Supplement material types not covered in self-construction data to enhance diversity.
3. High Quality: Pay attention to the quality of datasets during screening (whether uni-modal answers are possible, accuracy, reasonableness, etc.). For most world senses that can be answered with uni-modal data, they will not be included.
4. Discriminative: Pay attention to the performance of this data subset on the model during screening, and remove overly difficult subsets with little discriminative power.
3.3.2 Dataset Compression
Regarding the existing large-scale uni-modal benchmarks, to reduce the evaluation cost of large-scale models, we designed a clustering-guided hierarchical sampling method. This method achieves efficient compression and high evaluation consistency through the following steps:
Question Characterization: Represent each question as an x-dimensional vector, where dimensions correspond to scores from different models on that question.
Cluster-based Stratification: Utilize the K-Means algorithm to categorize questions into k clusters, each representing a "model performance similar" question type (e.g., easy questions, difficult questions, etc.).
Hierarchical Sampling: Determine the sample size for each stratum based on cluster size proportions, and construct the final evaluation subset through simple random sampling.
Validity Verification: Use Margin of Error (MoE) and Root Mean Square Error (RMSE) to measure theoretical and practical errors respectively. Verify the consistency between model rankings and scores using Spearman Rank Correlation Coefficient (SRCC) and Pearson Linear Correlation Coefficient (PLCC).
Finally, in order to ensure statistical stability, we repeat above steps by employing 5 random splits and 10-fold cross-validation. This method determines the optimal sample size through cost-benefit curve analysis, reducing evaluation costs by over 90% while maintaining accuracy (Figure.8).
3.4 Multi-Step Open-Ended Questions
Cross-modality problem requires the model to understand among multiple modalities with multi-step reasoning. However, existing evaluation benchmarks primarily focus on isolated problems where questions and answers remain independent of each other. Such a single-step evaluation paradigm cannot effectively assess the ability of a model to understand the complex, cross modality tasks that require multiple logical steps. Besides, most benchmarks adopt multiple-choice questions to verify the model’s performance, which reduces the difficulty of questions (e.g. allows for guessing) and may underestimate the model’s deficiencies in complex tasks. To address these issues, we introduce a new evaluation method, Multi-Step Open-Ended Questions (MO), which breaks down complex questions into multiple interdependent sub-tasks and requires the model to solve them step by step in the form of open-ended text. The construction process and example are shown in Figure.5. The multi-step open-ended question is worth 10 points. During the process of providing a reference answer, scores are assigned to each sub-question based on its difficulty and necessity for answering the final question, which facilitates the evaluation of the model’s accuracy at each step.
Experiments (4.4.1) show that compared with single-step evaluation method (e.g. multiple-choice questions), multi-step open-ended questions can effectively observe the ability decay of models in long-chain reasoning, providing a more realistic difficulty for advanced models with stronger discrimination.
Figure 5: Example of multi-step open-ended questions.
Table 2: Cross-modal benchmark results on multimodal models
Model Visual Audio Omni-MC Omni-MO
MiniCPM-o-2.6 41.85 63.10 26.77 21.92
Qwen-2.5-omni-3B 45.71 65.81 27.26 25.72
Baichuan-omni-1.5 49.25 66.50 30.32 20.48
Ming-lite-omni-1.5 48.68 71.70 32.10 26.08
Qwen-2.5-omni-7B 52.81 68.11 33.23 29.12
Gemini-2.0-Flash 66.80 74.41 46.26 40.96
Gemini-2.5-Flash 76.64 75.20 68.23 62.00
Gemini-2.5-Pro 80.10 78.70 76.61 68.84
Table 3: Cross-modal analysis on model ability system
Model Perception Reasoning Overall
Cross-modal
Alignment
Cross-modal
Recognition
Overall
Sequential
Reasoning
Spatial
Reasoning
General
Reasoning
Overall
MiniCPM-o-2.6 25.55 36.11 27.75 38.78 22.94 26.36 26.40 26.77
Qwen-2.5-omni-3B 27.74 50.00 32.37 38.78 27.52 22.09 25.28 27.26
Baichuan-omni-1.5 27.74 47.22 31.79 28.57 35.78 27.52 29.75 30.32
Ming-lite-omni-1.5 32.85 55.56 37.57 30.61 30.28 29.46 29.98 32.10
Qwen-2.5-omni-7B 40.88 52.78 43.35 38.78 30.28 27.13 29.31 33.23
Gemini-2.0-Flash 46.72 61.11 49.71 63.27 35.78 44.19 44.97 46.29
Gemini-2.5-Flash 67.88 77.78 69.94 85.71 41.28 75.19 67.56 68.23
Gemini-2.5-Pro 81.02 83.33 81.50 91.84 55.05 78.68 74.72 76.61
4 Experiment and Analysis
4.1 Experiment Setting
We evaluate omni-models that support text, visual, and audio inputs simultaneously, including open-source models: Qwen-3-omni(Xu et al., 2025b), Qwen-2.5-omni-3B, Qwen-2.5-omni-7B(Xu et al., 2025a), Baichuan-omni-1.5(Li et al., 2025), MiniCPM-o-2.6(Yao et al., 2024), and Ming-lite-omni-1.5(AI et al., 2025), as well as closed-source models: Gemini-2.5-Pro, Gemini-2.5-Flash, and Gemini-2.0-Flash(Comanici et al., 2025). All model integrations strictly adhere to official implementations. In video processing, each model receives raw video and performs frame sampling according to its own sampling strategy.
In the following sections, we conduct comprehensive experiments on MMAO-Bench and try to answer the following questions:
1. How about the performance of current omni-models and their shortage abilities?
2. What is the relationship of uni-modal ability and omni-modal ability?
3. Could the MMAO-Bench provide an effective "All in One" evaluation for omni-models?
4.2 Model Performance
4.2.1 Overall Analysis
As shown in Table 2, the proprietary Gemini series of models substantially outperform the evaluated open-source counterparts across all benchmarks. Gemini-2.5-Pro consistently leads, establishing a clear state-of-the-art.
When analyzing the cross-modal tasks specifically, we observe distinct patterns. On the Omni-MC (Multiple-choice) benchmark, the performance of smaller open-source models like MiniCPM-o-2.6 (26.77) and Qwen-2.5-omni-3B (27.26) approximates that of random guessing (0.25). In contrast, larger models such as Qwen-2.5-omni-7B (33.23) demonstrate nascent capabilities beyond this baseline. Nevertheless, a significant performance gap remains, with the top open-source model trailing Gemini-2.5-Pro (76.61) by over 43 points.
The Omni-MO (multi-step open-ended) benchmark, which demands more complex reasoning, proves to be even more challenging for all models. Performance universally drops compared to Omni-MC. Most open-source models score poorly, with the best among them, Qwen-2.5-omni-7B, achieving only 29.12. This indicates severe limitations in their ability to perform multi-step cross-modal reasoning. While the Gemini models also see a score reduction, they maintain a commanding lead, with Gemini-2.5-Pro (68.84) showcasing a far more robust, though not perfect, reasoning capability. This highlights that complex, multi-step, cross-modal understanding remains a critical and unresolved challenge, especially for the open-source community.
In terms of perception capabilities, cross-modal alignment is more challenging than cross-modal recognition. Among open-source models, Qwen-2.5-omni-7B has the strongest alignment capability at 40.88, while Ming-lite-omni-1.5 has the strongest recognition capability at 55.56. The Gemini series significantly outperforms other models, and Gemini-2.5-Pro shows no significant weakness in alignment compared to its recognition capability, both exceeding 80.
In terms of reasoning capabilities 3, spatial reasoning is currently the most challenging task, with the best-performing Gemini-2.5-Pro achieving only 55.05. Among open-source models, Baichuan-omni-1.5 performs the best with a score of 35.78. Ming-lite-omni-1.5 demonstrates the best performance in complex reasoning and general reasoning among open-source models.
Overall, perception ability is relatively simpler compared to reasoning ability, and the model gap in reasoning ability is more significant. Taking Qwen-2.5-omni-7B and Gemini-2.5-Pro as examples, they differ by 38.2 points in perception and 45.41 points in reasoning. This demonstrates that reasoning ability is an extension of perception ability.
4.2.2 Top-tier Analysis
Why Gemini-2.5-Pro performances significantly better than other models? We attempt to provide an analysis and some hypotheses. On one hand, it stems from the leading uni-modal understanding ability. On the other hand, from the perspective of the reasoning process, Gemini possesses audio captioning capabilities, and it can also naturally incorporate audio content as a basis for reasoning. Existing open-source models, due to their smaller size, lack reasoning processes in a multimodal context. Limited reasoning mostly relies on text or images, with few involving specific audio content.
However, does Gemini performance good enough? To answer this question, we invited human experts to conduct a competition.
Finding 1. Gemini-2.5-Pro has reached human comparable perception ability in omni-modal understanding, but still has a performance gap on reasoning ability. Compared to human experts, Gemini-2.5-Pro only has 8.3% less performance, which shows humanity comparable intelligence. The comparison of scores for specific ability items can be seen in Figure6. Dived into ability analysis, we find a interesting phenomena that human is more capable on reasoning than perception (81.3% > 74.3%), and perception ability is even lower than model’s. Through interviewing some of the human experts, we find that human may lost some information in video or audio in one hand, and in the other hand human has limited world knowledge than large language models.
Figure 6: The competition between Human experts and Gemini-2.5-Pro. Gemini-2.5-Pro shows comparable perception capability but lower reasoning capability.
4.3 Uni-Modal v.s. Cross-Modal
To investigate the relationship of uni-modal and omni-modal understanding ability, we conduct regression analysis and ablation experiments. Thanks to the unified ability systems and the high quality of omni-modal samples in MMAO-Bench, we find some interesting observations.
Finding 2. Compositional Law: the performance of omni-modal ability is power-law related to the multiply of uni-modals’ performances. Observing the Table.2, we find that the omni-modal performances are correlation to the multiply of their uni-modals’ performances. In details, the performances of different models are relatively close in audio understanding but show significant differences in vision understanding, and these differences are also pronounced in the omni-modal considering both abilities. Gemini-2.5-Flash and Gemini-2.0-Flash are similar in audio understanding, but for vision modality, Gemini-2.5-Flash outperforms Gemini-2.0-Flash by 9.8, and it is 24.4 stronger in the cross-modal understanding.
Finding 3. The weaker models show "short-board effect", while the stronger models demonstrate emergent ability. We draw the multiply of uni-modal scores as a dot line in the Figure.7. Compared with the cross-modal scores, models around the "random guessing stage" are below the line, while Gemini-2.5-Pro shows a significant gap above the line. Besides, these models demonstrate a power-law tendency that the stronger uni-modal ability can stimulate to much stronger cross-modal ability. This also emphasis that multiple modalities can enhance each other and lead to a higher intelligence.
Next, we conduct ablation studies to dig dive about the enhancement from vision and audio modality respectively.
Figure 7: Relationship between uni-modal ability and cross-modal ability among omni-models.
Table 4: Ablation of visual understanding ability.
Model Perception Reasoning Overall
Audio +Caption +Visual Audio +Caption +Visual Audio +Visual
MiniCPM-o-2.6 26.01 30.06 27.75 24.83 25.28 26.40 25.16 26.77
Qwen-2.5-omni-3B 21.39 29.48 32.37 22.60 24.83 25.28 22.26 27.26
Baichuan-omni-1.5 22.54 30.06 31.79 28.86 31.54 29.75 27.10 30.32
Ming-lite-omni-1.5 26.59 27.17 37.57 25.06 24.38 29.98 25.48 32.10
Qwen-2.5-omni-7B 26.01 33.53 43.35 20.81 25.28 29.31 22.26 33.23
Gemini-2.0-Flash 31.21 53.18 49.71 34.90 51.01 44.97 33.87 46.29
Gemini-2.5-Flash 33.53 68.21 69.94 44.30 70.02 67.56 41.29 68.23
Gemini-2.5-Pro 41.62 78.03 81.50 50.34 73.60 74.72 47.90 76.61
4.3.1 Ablation Visual Understanding
To evaluate the enhancement of visual information on model understanding, we sequentially assessed the performance of Audio (only auditory information), +Caption (auditory + Caption information), and +Visual (auditory information + visual information), where captions were derived from descriptions of videos by Gemini-2.5-Pro.
In the case of audio-only input, most models’ metrics are significantly lower than when both auditory and visual information are provided, and the metrics exhibit a random distribution around 25%. Gemini-2.5-Pro identified problem shortcuts through methods such as elimination, achieving a score of 47.9, which will be optimized in subsequent versions.
After providing the caption information, all model metrics showed significant increases, but the magnitude of the rise varied considerably. In terms of perceptual ability, the Gemini series generally improved by 20 points, while the Qwen-2.5-omni series and Baichuan-omni-1.5 saw an improvement of over 8 points. The MiniCPM-o-2.6 showed a smaller increase of around 4 points, whereas the Ming-lite-omni-1.5 barely improved at all. This indicates that different models have varying abilities to reconstruct image content based on textual descriptions, and similar phenomena are observed in reasoning capabilities as well.
Compared with caption information, the original visual input brings more information. On models with better baseline capabilities, such as Ming-lite-omni-1.5 and Qwen-2.5-omni-7B, caption information is far less than the original visual information. On models with slightly weaker baseline capabilities, the two are approximately the same. Interestingly, the Gemini series has a very high utilization rate of caption information, and even Gemini-2.0-Flash performs better with +caption than with +visual.
4.3.2 Ablation Audio Understanding
We categorize sound types into ambient sounds and dialogue, and provide their transcription versions based on Gemini-2.5-pro, where dialogue is given dialogue text, and ambient sounds are given sound descriptions.
Similar to visual information, in the case of only providing auditory information, most models’ metrics are significantly lower than when both auditory and visual information are provided; after providing speech transcripts, the improvement is again more pronounced in the Gemini series.
MiniCPM-o-2.6 performs better than +Audio on +ASR, confirming that the bottleneck of this model lies in audio understanding.
Table 5: Ablation of audio understanding ability.
Model Perception Reasoning Overall
Video +ASR +Audio Video +ASR +Audio Video +ASR +Audio
MiniCPM-o-2.6 27.75 31.79 27.75 27.07 29.98 26.40 27.26 30.48 26.77
Qwen-2.5-omni-3B 27.17 30.06 32.37 18.79 23.94 25.28 21.13 25.65 27.26
Baichuan-omni-1.5 26.01 27.17 31.79 27.07 23.04 29.75 26.77 24.19 30.32
Ming-lite-omni-1.5 32.37 35.84 37.57 23.04 27.52 29.98 25.65 29.84 32.10
Qwen-2.5-omni-7B 30.06 41.04 43.35 24.16 27.52 29.31 25.81 31.29 33.23
Gemini-2.0-Flash 31.79 55.49 49.71 29.08 47.43 44.97 29.84 49.68 46.29
Gemini-2.5-Flash 38.73 72.83 69.94 37.58 68.68 67.56 37.90 69.84 68.23
Gemini-2.5-Pro 46.82 76.88 81.50 47.20 75.62 74.72 47.10 75.97 76.61
In terms of environmental sound, most open-source models perform poorly. The two more capable models, Ming-lite-omni-1.5 and Qwen-2.5-omni-7B, show better performance in dialogue than in environmental sound. Additionally, we observed that the audio information of MiniCPM-o-2.6 and Ming-lite-omni-1.5 has side effects in environmental sound scenarios. When audio descriptions replace the original audio, both models show significant improvement, indicating that Ming-lite-omni-1.5 and Qwen-2.5-omni-7B have relatively weaker understanding capabilities of audio elements.
In conversational scenarios, the Gemini series achieves ASR results that are nearly on par with, or even surpass, the original audio quality. This distinctive feature is absent in most other open-source models, with the exception of MiniCPM-o-2.6, where the +ASR results exceed +Audio 4 points.
Table 6: Ablation of audio understanding ability.
Model Environment Speech Overall
Visual +Caption +Audio Visual +ASR +Audio Visual +ASR/Caption +Audio
MiniCPM-o-2.6 33.33 31.11 28.89 26.70 30.54 26.53 27.26 30.48 26.77
Qwen-2.5-omni-3B 15.56 31.11 28.89 21.64 25.31 27.23 21.13 25.65 27.26
Baichuan-omni-1.5 26.67 22.22 33.33 26.70 24.43 30.02 26.77 24.19 30.32
Ming-lite-omni-1.5 35.56 37.78 28.89 24.96 29.32 32.29 25.65 29.84 32.10
Qwen-2.5-omni-7B 28.89 40.00 26.67 25.65 30.72 33.86 25.81 31.29 33.23
Gemini-2.0-Flash 37.78 42.22 40.00 29.32 50.44 46.77 29.84 49.68 46.29
Gemini-2.5-Flash 53.33 62.22 51.11 36.47 70.51 69.63 37.90 69.84 68.23
Gemini-2.5-Pro 57.78 71.11 75.56 46.07 76.44 76.79 47.10 75.97 76.61
4.4 Benchmark Analysis
In this section, we verify the effectiveness of MMAO-Bench on three aspects, the performance of multi-step open-ended question, the performance of dataset compression and the benchmark comparison with other open-sourced benchmarks.
4.4.1 Multi-Step Open-Ended Question Analysis
In this work, we introduce a new type of evaluation method, multi-step open-ended question, which effectively assess the complex reasoning ability, especially appears in cross-modality understanding. As shown in Table.7, the experimental results demonstrate the stratified performance among different models. Gemini-2.5-Pro leads with a comprehensive score of 68.84, showcasing its outstanding capabilities as a top-tier closed-source model. Following closely are Gemini-2.5-Flash and Gemini-2.0-Flash, which also form a clear leading tier. Among open-source models, Qwen-2.5-omni-7B performs best, significantly outperforming other smaller-scale open-source models. This indicates that model scale and advanced architecture remain key factors in addressing complex multi-question tasks.
As the depth of questions increases, model performance generally declines. This is a very significant trend, also validating the effectiveness of our dataset design. For almost all models, their performance on small question 1 is better than on small question 2, while the score for small question 3+ is the lowest. For example, Gemini-2.5-Pro’s score on the "perception" task dropped from 26.60 on small question 1 to 21.79 on small question 3+; on the "reasoning" task, the score sharply decreased from 38.2 on small question 1 to 21.33 on small question 3+. This performance decay phenomenon reveals the common challenges current large models face in handling long-range dependencies, maintaining conversational context, and conducting multi-step chain-of-thought reasoning. Initial perception and simple reasoning are relatively easy, but as conversations deepen, the accumulation of errors and loss of context lead to significantly lower success rates in answering more complex subsequent questions.
Reasoning ability is the key bottleneck that distinguishes model performance. Comparing the total scores of "perception" and "reasoning" dimensions, we found that all models scored higher on perception tasks than on reasoning tasks. For example, Gemini-2.5-Pro’s perception total score is 78.4, while its reasoning total score is 66.45. Among open-source models, Qwen-2.5-omni-7B’s perception total score (40.8) is significantly higher than its reasoning total score (26.20). This indicates that although models have already acquired basic perception capabilities (such as object recognition and sound recognition), there is still significant room for improvement in their ability to transform this perceptual information into complex logical, spatial, or causal reasoning. Our dataset successfully amplifies this capability gap through its progressive question design.
Table 7: Performance on Multi-Step Open-Ended Questions.
Model Perception Reasoning Overall
Q1 Q2 Q3+ All Q1 Q2 Q3+ All Q1 Q2 Q3+ All
Baichuan-omni-3B 8.95 7.75 5.38 24.2 11.8 9.6 9.33 19.55 9.52 8.12 5.87 20.48
MiniCPM-6-2.6B 10.65 9.35 8.11 34.2 22.4 11.8 6.67 18.85 11.16 8.12 6.33 21.92
Qwen-2.5-omni-3B 10.95 8.59 8.76 34.2 19.4 13.2 5.33 23.60 12.64 9.52 8.33 25.72
Ming-lite-omni 8.35 7.20 6.29 33.2 20.8 12.4 4.67 24.30 12.68 9.96 7.69 26.08
Qwen-2.5-omni-7B 12.85 9.75 7.90 40.8 18.4 17.6 16.00 26.20 13.96 11.33 8.92 29.12
Gemini-2.0-Flash 17.55 13.60 11.32 56.2 27.8 22.0 21.33 37.15 19.6 15.28 12.56 40.96
Gemini-2.5-Flash 23.40 25.05 20.66 72.4 35.2 31.4 26.00 59.40 25.76 26.32 21.32 62.00
Gemini-2.5-Pro 26.60 29.10 21.79 78.4 38.2 33.8 21.33 66.45 28.92 30.04 21.74 68.84
4.4.2 Dataset Compression
We design a cluster-guided stratified sampling to compress the scale of benchmark. To evaluate the consistency of model ranking and the best size of compression data size, we conduct several experiments to analysis.
Baseline dataset consists of  8000 samples including 15 open-source benchmarks (e.g. MathVista and MMAU, details see appendix) and 12 models evaluation results on them, which splits into 8/4 on models as training/test set. Kmeans is used to cluster with K=48. To eliminate the random factor, we conduct 5-fold settings and evaluate 10 times on each setting.
To verify the compression performance, we use metrics Margin of Error(MoE), Root Mean Square Error(RMSE), CI Coverage, Spearman’s linear correlation coefficient(SRCC) and Pearson’s linear correlation coefficient(PLCC). The experimental result is shown in Figure.8, when compression rate is 10%, the coverage is nearly 100% and PLCC is above 0.98, SRCC is above 0.98, RMSE is 0.06, which means the model ranking is consistent.
Figure 8: Data compression Performance.
4.4.3 Benchmark Comparison
Besides dataset quality and evaluation efficiency, a good benchmark should provide sufficient metric discriminability and a progressive difficulty scale to drive model capability growth. Compared with existing omni-modal benchmarks (including OmniBench(Li et al., 2024b), WorldSense(Hong et al., 2025) and AVOdyssey(Gong et al., 2024)), MMAO-Bench has relatively consistent rankings but a higher standard deviation of 19.5 in evaluation scores, while the best other OmniBench only has 12.75, showing the MMAO-Bench has greater discriminative ability. A simple way to improve discrimination is increasing problem difficulty, which can’t guide model’s growth. As a verification, we can see that Gemini-2.5-pro scored 76.61 on MMAO-Bench, but only 65.49 on OmniBench, indicating that MMAO-Bench has appropriate difficulty levels.
Figure 9: Comparison between omni benchmarks.
5 Conclusion
In this work, we introduce a high quality and diversity benchmark to evaluate omni models comprehensively. The main foundings are summarized as below.
1. With unified data framework in MMAO-Bench, we found that the multimodal capability may not simply be a linear superposition of uni-modalities, but rather follows a significant multiplicative relationship. The evaluation results show that it manifests as a bottleneck effect on weak models, while exhibiting synergistic promotion on strong models.
2. Both uni-modal and omni-modal understanding capability of the Gemini series far surpasses existing open-source omni-models. The Gemini-2.5-Pro shows comparable perception capability with human experts but still has a performance gap in reasoning aspect.
3. Besides better dataset quality and evaluation efficiency, MMAO-Bench can provide sufficient metric discriminability and a progressive difficulty scale to drive model capability growth.
In the future work, we will extend the dataset’s scale by human-in-the-loop automatic pipeline and hold a private test set to avoid hacking. The ability coverage also needs to extend to more difficult reasoning tasks like STEM and code. At the same time, the relationship among cross-modals understandings and how to improve them are still exciting problem to explore.

### v2

@@
-MMAO-Bench: MultiModal All in One Benchmark Reveals Compositional Law between Uni-modal and Omni-modal in OmniModels
+UNO-Bench: A Unified Benchmark for Exploring the Compositional Law Between Uni-modal and Omni-modal in OmniModels
 Abstract
@@
 However, the correlation between uni-modal and omni-modal remains unclear, which requires comprehensive evaluation to drive omni model’s intelligence evolution.
-In this work, we propose a novel, high quality and diversity omni model benchmark, MultiModal All in One Benchmark (MMAO-Bench), which effectively assesses both uni-modal and omni-modal understanding capabilities.
-The benchmark consists of 1880 human curated samples, across 44 task types, and a innovative multi-step open-ended question type that better assess complex reasoning tasks.
-Experimental result shows the compositional law between cross-modal and uni-modal performance and the omni-modal capability manifests as a bottleneck effect on weak models, while exhibiting synergistic promotion on strong models.
-MMAO-Bench: MultiModal All in One Benchmark Reveals Compositional Law between Uni-modal and Omni-modal in OmniModels
+In this work, we propose a novel, high quality and UNified Omni model benchmark, UNO-Bench, which effectively assesses both UNi-modal and Omni-modal capabilities.
+The benchmark consists of 3730 human curated samples, with 98% cross-modality solvability, across 44 task types, and an innovative multi-step open-ended question type for assessing complex reasoning.
+Besides, a general scoring model supporting 6 question types is proposed for automated evaluation with 95% accuracy.
+Experimental result shows the Compositional Law between omni-modal and uni-modal performance and the omni-modal capability manifests as a bottleneck effect on weak models, while exhibiting synergistic promotion on strong models.
+Our code and data are available at GitHub.
+UNO-Bench: A Unified Benchmark for Exploring the Compositional Law Between Uni-modal and Omni-modal in OmniModels
 Chen Chenthanks: Equal contribution.
@@
 The evaluation paradigm for these models has consequently expanded from assessing uni-modal understanding capabilities (i.e.
-visual understanding, audio understanding) to the next-level of intelligence (i.e.
-omni-modal understanding).
-Existing omni-model evaluation benchmarks remain relatively scarce and focus on different aspects.
+visual understanding, audio understanding) to the next-level of intelligence, omni-modal understanding.
+Existing omni model evaluation benchmarks remain relatively scarce and focus on different aspects.
 For example, some prioritize image comprehension(Li et al., 2024b), others emphasize video understanding(Hong et al., 2025), while a subset concentrates on speech interaction(Gong et al., 2024).
 Notably, existing datasets are exclusively English-centric, lacking evaluation benchmarks for Chinese linguistic contexts.
-The ideal omni-model should simultaneously preserve visual understanding capabilities (e.g., MMBench(Liu et al., 2024a)/MathVista(Lu et al., 2024)/MVBench(Li et al., 2024a)), speech interaction proficiency (e.g., MMAU(Sakshi et al., 2025)), and cross-modal integration capacity (e.g., OmniBench(Li et al., 2024b)/WorldSense(Hong et al., 2025)).
+The ideal omni model should simultaneously preserve visual understanding capabilities (e.g., MMBench(Liu et al., 2024b)/MathVista(Lu et al., 2024a)/MVBench(Li et al., 2024a)), speech interaction proficiency (e.g., MMAU(Sakshi et al., 2025)), and cross-modal integration capacity (e.g., OmniBench(Li et al., 2024b)/WorldSense(Hong et al., 2025)).
 Current evaluation paradigms employ disjointed benchmark suites for separate capability testing, creating resource-intensive evaluation processes and disconnected modality assessments.
 Beyond uni-modal, omni-modal capability introduces advanced challenges across image, video and audio modality.
-However, 77% questions from WorldSense are solvable without video or audio, and 25% questions from OmniBench contain erroneous answers.
-These designs limit the evaluation and analysis of omni models’ abilities.
-Addressing theses limitations, we propose a novel, high quality and diversity omni model benchmark, MultiModal All in One Benchmark (MMAO-Bench), that enables comprehensive model assessment and pushing omni model to the next-level of intelligence.
+However, 77% questions from WorldSense are solvable without vision or audio, and 25% questions from OmniBench contain erroneous answers.
+These issues limit the evaluation and analysis of omni models’ capabilities.
+Dataset Omni-modal Uni-modal Acc. Solvable Source #Tasks #QA Pairs QA Type Language
+MMBench I - - 80% private 20 3217 MC EN/CH
+MMAU A - - 15% private 27 10000 MC EN
+MVBench V - - public 20 4000 MC EN
+OmniBench I+A 75% 90% public 8 1142 MC EN
+AV-Odyssey I+V+A 91% 99% public 26 4555 MC EN
+WorldSense V+A 99% 23% public 26 3172 MC EN
+Daily-Omni V+A 94% 59% public 6 1197 MC EN
+UNO-Bench-omni I+V+A -  100% 98% 90% private 44 1250 MC/MO EN/CH
+UNO-Bench-uni - I/V/A 86% 97% 42% private 44 2480 MC EN/CH
+Table 1: Comparison of MultiModal Benchmarks, with I, A, V, and T representing image, audio, video, and text modalities, respectively.
+It reports on the accuracy of question-answer pairs and the percentage of questions requiring omni-modal solutions, labeled as Acc.
+and Solvable.
+The Source category specifies the origin of the materials.
+Private sources, as opposed to public ones, can prevent data contamination.
+QA types include MC for multi-choice questions and MO for multi-step open-ended questions.
+EN and CH denote English and Chinese languages.
+UNO-Bench includes 1250 human-curated samples in the omni-modal section (referred to as -omni) and 2480 enhanced samples in the uni-modal section (referred to as -uni).
+Due to the limited quality and coverage of existing benchmarks, we introduce a novel and unified benchmark UNO-Bench.
+As shown in Figure.1, the materials are collected from human crafting which prevents data contamination while better aligning with real-world scenarios.
+Beyond conventional multiple-choice questions, the evaluation adopts an innovative Multi-Step Open-Ended Question type to show a more realistic and discriminative evaluation result on complex reasoning.
+Besides the human crafted dataset, we incorporate existing uni-modal datasets by aggregating them systematically and design a clustering-guided sampling method to achieves both evaluation efficiency and consistency.
+In this way, our benchmark involves a comprehensive assessment necessitating omni models to maintain their uni-modal ability while simultaneously acquiring omni-modal capability.
 Main Contributions:
 1.
-Propose the first MultiModal All in One Benchmark evaluation framework specifically designed for omni models, which efficiently assesses both uni-modal and omni-modal understanding capabilities.
-This framework verifies the compositional law between omni-modal and uni-modal capability.
-For weaker models, this manifests itself as a "short-board effect" (where performance is constrained by the weakest modality), while stronger models demonstrate emergent ability across modalities.
+Propose the first UNified Omni model benchmark, UNO-Bench, which efficiently assesses both UNi-modal and Omni-modal understanding capabilities.
+UNO-Bench verifies the Compositional Law between omni-modal and uni-modal capability.
+The omni-modal capability acts as a bottleneck effect on weaker models, but shows synergistic enhancement on stronger models.
 2.
-Establish a systematic and unified dataset framework, featuring complete manual annotation to ensure high quality and diversity.
-This approach effectively prevents data contamination while better aligning with practical application scenarios.
-As a result, the MMAO-Bench comprises 44 different types of tasks and is 100% passing the quality check, which significantly surpassing existing datasets.
+Establish a high quality and diversity dataset construction pipeline including human-centric process and automated data compression.
+As as result, UNO-Bench comprises 1250 human curated samples for omni-modal with 98% cross-modality solvability, and 2480 enhanced samples for uni-modal, across 44 task types and 6 question types.
+The human-created novel dataset is well-suited to real-world scenarios, particularly within the Chinese context, whereas the automatically compressed dataset offers a 90% increase in speed and maintains 98% consistency across 18 public benchmarks.
+Its comprehensive quality and efficiency significantly surpasses existing datasets.
 3.
-Beyond conventional multiple-choice formats, the evaluation incorporates innovative Multi-Step Open-Ended Question tasks to show a more realistic and discriminative evaluation result on complex reasoning.
+Beyond conventional multiple-choice question type, the evaluation incorporates innovative Multi-Step Open-Ended Question (MO) to show a more realistic and discriminative evaluation result on complex reasoning especially for multi-step reasoning across modalities.
+For automated evaluation, a General Scoring Model is proposed to support 6 kinds of question types with 95% accuracy on OOD models and benchmarks.
+Figure 1: Illustration of the unified ability taxonomy proposed in UNO-Bench.
+(a) exemplarily presents the way to unify uni-modal and omni-modal samples.
+(b) shows the three-layer hierarchical structure of the unified ability taxonomy and its distribution in UNO-Bench.
+Among the 13 abilities, (c) demonstrates the $4$ diverse task types that cover a large range of real world scenarios.
 2 Related Work
@@
 Based on large language models, vision language models (VLM) (Bai et al., 2025; Xiaomi, 2025; Zeng et al., 2025) and audio language models (ALM) (Ding et al., 2025; Wu et al., 2025) introduce the general intelligence to vision modality and audio modality respectively.
-Various uni-modal benchmarks conduct comprehensive evaluations on VLMs (Liu et al., 2024a; Lu et al., 2024; Wang et al., 2024a, b; Liu et al., 2024c; Mathew et al., 2021; Ouyang et al., 2024; Li et al., 2024a; Wu et al., 2024; Liu et al., 2024b; xAI, 2023; Xiao et al., 2021; Huang et al., 2025; Hu et al., 2025; Fu et al., 2024) and ALMs (Ardila et al., 2019; Wang et al., 2021; Yang et al., 2024; Ao et al., 2024).
-For image modality, MMBench(Liu et al., 2024a) proposed a systematically designed benchmark to evaluate general image understanding on 20 different tasks.
-Focused on mathematic, MathVision(Wang et al., 2024a) collected questions from 19 mathematic competitions to evaluate VLMs complex reasoning ability.
-In addition to above, OCRBench(Liu et al., 2024c) supplied the evaluation on text recognition and document understanding.
+Various uni-modal benchmarks conduct comprehensive evaluations on VLMs (Liu et al., 2024b; Lu et al., 2024a; Wang et al., 2024b, c; Liu et al., 2024d; Mathew et al., 2021; Ouyang et al., 2024; Li et al., 2024a; Wu et al., 2024; Liu et al., 2024c; xAI, 2023; Xiao et al., 2021; Huang et al., 2025; Hu et al., 2025; Fu et al., 2024) and ALMs (Ardila et al., 2019; Wang et al., 2021; Yang et al., 2024; Ao et al., 2024).
+For image modality, MMBench(Liu et al., 2024b) proposed a systematically designed benchmark to evaluate general image understanding on 20 different tasks.
+Focused on mathematic, MathVision(Wang et al., 2024b) collected questions from 19 mathematic competitions to evaluate VLMs complex reasoning ability.
+In addition to above, OCRBench(Liu et al., 2024d) supplied the evaluation on text recognition and document understanding.
 For video modality, MVBench(Li et al., 2024a) aggregated 11 public video benchmarks and incorporated data enhancement process to cover 20 dynamic video understanding tasks.
@@
 2.2 Omni-Modal Benchmarks
-Omni models have arisen in recent years(Comanici et al., 2025; Xu et al., 2025b; AI et al., 2025; Li et al., 2025), as the pioneer, Gemini(Comanici et al., 2025) shows a strong ability in understanding both vision and audio, while qwen-3-omni(Xu et al., 2025b) provides leading performance in open-sourced models.
+Omni models have arisen in recent years(Comanici et al., 2025; Xu et al., 2025b; AI et al., 2025; Li et al., 2025), as the pioneer, Gemini(Comanici et al., 2025) shows a strong ability in understanding both vision and audio, while Qwen-3-Omni(Xu et al., 2025b) provides leading performance in open-source models.
 However, there’s less omni-modal benchmarks that can evaluate the modality combination across image, video and audio.
@@
 Other datasets focus on audio (Gong et al., 2024) or video (Zhou et al., 2025) and cover limited task types.
-Due to the limited quality and coverage of existing benchmarks, we introduce a new benchmark MMAO-Bench which comprise 1250 human crafted questions for omni-modal and 630 enhanced samples for uni-modal, across 44 task types with 98% cross-modality solvable.
-The materials are collected from in-house dataset with manual annotation which prevents data contamination while better aligning with practical application scenarios.
-Beyond conventional multiple-choice formats, the evaluation incorporates innovative multi-step open-ended question tasks to show a more realistic and discriminative evaluation result on complex reasoning.
-Besides the self-constructed omni-modal dataset, we construct a uni-modal dataset by aggregating public uni-model datasets systematically and design a clustering-guided sampling method to achieves both evaluation efficiency and consistency.
-In this way, our benchmark can have a thoroughly evaluation that requires omni models to keep uni-modal abilities and obtain omni-modal abilities at the same time.
-Dataset Omni-modal Uni-modal Acc. Solvable Source #Tasks #QA Pairs QA Type Language
-MMBench I - - 80%private 20 3217 MC EN/ZH
-MMAU A - - 15%private 27 10000 MC EN
-MVBench V - - public 20 4000 MC EN
-OmniBench I+A 75% 90% public 8 1142 MC EN
-AV-Odyssey I+V+A 91% 99% public 26 4555 MC EN
-WorldSense V+A 99% 23% public 26 3172 MC EN
-Daily-Omni V+A - - public 6 1197 MC EN
-MMAO-Bench I+V+A I/V/A 100% 98% 90%private 44 1880 MC/MO EN/ZH
-Table 1: Comparison of MultiModal Benchmarks.
-Where I,A,V and T indicate image, audio, video and text modality.
-The accuracy of question answer pairs and the percent of questions needing omni-modal to solve is shown in Acc.
-and Solvable.
-Source indicates where the materials are collected from, public datasets or private datasets that can prevent data contamination.
-For QA type, MC, SO and MO indicates multi-choise QA, single-open question QA and multi-open question QA.
-EN and ZH indicates english and chinese.
+Addressing these limitations, we propose a novel and unified benchmark, UNO-Bench, that enables comprehensive model assessment and pushing omni model to the next-level of intelligence.
 3 Method
-In this section, we first introduce the comprehensive design of model ability system in Section.3.1.
-The the high-quality data construction pipeline are introduced in Section.3.2.
-The multi-step open-ended questions are introduced in Section.3.4.
-Finally, to improve the evaluation efficiency, we introduce a general evaluation dataset compression method in Section.3.3.2.
-3.1 Model Ability System
-As shown in Figure.1, the model’s capabilities are systematically categorized into two primary dimensions: Perception and Reasoning.
-Perception Layer Structured through six recognition types: Object Perception, Attribute Perception, Scenario Perception, Spatial Perception, Cross- Modal Conversion, Semantic Understanding.
+In this section, we first introduce the omni-modal dataset construction pipeline in Section.3.1.
+The uni-modal dataset improvement method and a general evaluation dataset compression method to improve the evaluation efficiency are introduced in Section.3.2.
+Finally, the multi-step open-ended questions are introduced in Section.3.3.
+3.1 Omni-modal Dataset Construction
+We have established a human-centric data construction pipeline (Figure.2) that efficiently empower human intelligence to produce high-quality and high-diversity dataset.
+Figure 2: Dataset Construction Pipeline includes human crafted process (left side) and automated data compression (right side).
+First, we collect diverse and novel materials to prevent data contamination.
+Second, with the proposed unified ability taxonomy, human annotators including experts will craft questions, answers and record audios in real-world scenarios.
+Finally, with model checking, ablation study and human experts revision, we achieves high quality and diversity dataset.
+Regarding automated data compression, we use model performances as the sample feature, employing clustering and resampling to achieve a 90% dataset compression.
+3.1.1 Model Ability Taxonomy
+Through cumulative experiences on multimodal evaluation from both model-side and user-side, we summarize the capabilities of uni-modal and omni-modal into a unified model ability taxonomy.
+As shown in Figure.1, the omni model’s capabilities are systematically categorized into two primary dimensions: Perception and Reasoning.
+Detailed definitions and examples can be found in the Appendix.E.
+Perception dimension structured through seven recognition types: Object Perception, Attribute Perception, Scenario Perception, Spatial Perception, Cross-Modal Conversion, Semantic Understanding.
 In addition, we incorporate Cross-Modal Alignment evaluation to assess information synchronization across modalities.
-Reasoning Layer extends conventional reasoning categories (General, STEM, Code) with: Spatial Reasoning, Temporal Reasoning, and Complex Reasoning (multi-conditional, multi-step problem solving).
-This ability system combines uni-modal and cross-modal abilities which enables cross-modal measurement that is particularly critical for advanced omni-models.
-Based on this system, we create a diversity dataset with 44 task types distributed as shown in Figure.2.
-Figure 1: Illustration of the omni-modal ability system proposed in MMAO-Bench.
-Figure 2: Data distribution in model abilities and corresponding task types.
-3.2 Omni-modal Dataset Construction
-We have established a standardized and efficient data construction process (Figure.3) that produces high-quality and high-diversity dataset.
-Figure 3: Data Construction Pipeline
-Preparation: (1) Task Design.
-Deconstruct cross-modal capability systems and task types.
-Details can be seen in Figure.2.
-(2) Material Collection.
+The perception dimension fuses visual and audio modalities. For example, Scenario Perception includes the recognition of visual scenes and the judgment of audio scenes.
+Reasoning dimension extends conventional reasoning categories (General, STEM, Code) with: Spatial Reasoning (including Static Reasoning and Dynamic Reasoning), Temporal Reasoning, and Complex Reasoning (multi-conditional, multi-step problem solving).
+The unified ability taxonomy combines uni-modal and omni-modal abilities which provides a comprehensive measurement that is particularly critical for advanced omni models.
+Based on this taxonomy, we create a diversity dataset with 44 task types distributed as shown in Figure.1.
+3.1.2 Material Collection
+In both data quality checks and experimental results, we found that natural audio-visual synchronized video data contains a large amount of information redundancy, only a few videos require both audio and vision modality simultaneously.
+Therefore, we begin with carefully designed material collection.
 Our materials have the following three characteristics: a.
 Diverse Sources.
-The majority of our materials are real-world photos and videos collected through crowdsourcing, with another portion sourced from copyright-free website.
-Additionally, a small fraction comes from high-quality public datasets such as MMVU, LongVideoBench, and VideoVista.
+The majority of our materials are real-world photos and videos collected through crowdsourcing, and another portion sourced from copyright-free website.
+Additionally, a small fraction comes from high-quality public datasets such as MMVU(Zhao et al., 2025), LongVideoBench(Wu et al., 2024), and VideoVista(Chen et al., 2025).
 b.
@@
 With over 20 participants in the recording process, the audio features are rich and closely reflect the diverse vocal characteristics of the real world.
-(3) Material filtering.
+Finally, we conduct material filtering.
 Eliminate meaningless, illogical, and low-quality materials, and categorize the remaining materials by theme to create a material library.
 Additionally, label the materials with more detailed information such as subject, event, scene, and style to facilitate subsequent annotators to quickly find matching materials.
-QA Annotation: (1) Annotators.
-It mainly includes two parts: human experts and high-quality crowd-sourced users.
+3.1.3 QA Annotation
+Our annotators consist of human experts and high-quality crowd-sourced users.
 Human experts have extensive experience in cross-modal data construction and annotation, a deeper understanding of model capabilities, and thus ensure higher professionalism and specificity in the data they construct.
 Most crowd-sourced users are college students with rich experience in multi-modal model interactions and diverse professional backgrounds, providing data with better authenticity and diversity.
-(2) Annotation process.
 First, annotators clarify the required image/video features based on task type definitions and filter appropriate materials from existing libraries using tags.
@@
 Through this workflow, we ultimately generate complete QA pairs encompassing three modalities: visual, auditory, and textual.
-Quality Inspection: We have established a multi-stage, cyclically validated quality assurance system composed of automated tools and manual review.
+Compared to conventional methods limited to human intervention only during the quality assurance phase, our pipeline integrates a "Human-in-the-Loop" approach, ensuring continuous manual involvement from the initial data sourcing to the final output.
+This methodology not only prevents data leakage but also more accurately simulates real-world scenarios.
+Furthermore, the manually curated Chinese dataset genuinely captures user requirements in a Chinese linguistic context, compensating for the shortcomings of most existing English-centric datasets.
+3.1.4 Quality Inspection
+To ensure the data quality, we have established a multi-stage, cyclically validated quality assurance system composed of automated tools and manual review.
 Each question undergoes at least three rounds of independent quality inspection to maximize data quality.
 First, a preliminary model check is conducted to filter out cases with ambiguous questions, non-unique answers, or those that do not conform to the task type.
-Secondly, through modality ablation experiments, we remove one modality of information from the QA pair to see if the model can answer based solely on the remaining information.
-If the question becomes unanswerable or ambiguous after removing any one modality, it proves the three-modality solvability of the data.
+Second, through modality ablation experiments, we remove one modality of information from the QA pair to see if the model can answer based solely on the remaining information.
+If the question becomes unsolvable or ambiguous after removing any one modality, it proves the cross-modality solvability of the data.
 Finally, manual quality inspection and revision are performed.
@@
 3.2.1 Quality Improvement
-Existing public uni-modal datasets are bothered by data leakage issue.
-To verify the influence, we adopt privatization improvement on the widely used public dataset MMBench.
-As shown in Figure4, experiments show that the performance of models have better distinguishability after dataset improvement, reflecting the true capability differences between models.
+Existing public uni-modal datasets are bothered by data leakage issue(Xu et al., 2024).
+To verify the influence, we adopt privatization improvement on the widely used public dataset MMBench(Liu et al., 2024b).
+As shown in Figure.3, the performance of models have better distinguishability after dataset improvement, reflecting the true capability differences between models.
 Therefore, for uni-modal data, we also follow the aforementioned construction process for self-construction datasets.
-Figure 4: Comparison result of privatization improvement. After improvement, the performances among models are more progressively and distinguishable.
+Figure 3: Comparison result of privatization improvement. After improvement, the performances among models are more distinguishable.
 In addition to self-constructed data, we also selected some multimodal data from public datasets to supplement in terms of capability items and data types.
-(Data mainly comes from AV-Odyssey and WorldSense, accounting for 11% of the total).
+(Data mainly comes from AV-Odyssey(Gong et al., 2024) and WorldSense(Hong et al., 2025), accounting for 11% of the total).
 The specific selection logic is as follows:
-1.
-Comprehensiveness: In terms of capabilities, focus on supplementing the perception part with a relatively low self-construction proportion, while also adding some reasoning questions; in terms of data types, prioritize selecting video + audio modality combinations with a lower self-construction proportion for supplementation, followed by image + audio.
-2. Diversity: Supplement material types not covered in self-construction data to enhance diversity.
-3.
-High Quality: Pay attention to the quality of datasets during screening (whether uni-modal answers are possible, accuracy, reasonableness, etc.).
-For most world senses that can be answered with uni-modal data, they will not be included.
-4. Discriminative: Pay attention to the performance of this data subset on the model during screening, and remove overly difficult subsets with little discriminative power.
+Comprehensiveness: In terms of capabilities, focus on supplementing the perception part with a relatively low self-construction proportion, while also adding some reasoning questions; in terms of data types, prioritize selecting the video plus audio modality combination with a lower self-construction proportion for supplementation, followed by image plus audio.
+Diversity: Supplement material types not covered in self-construction data to enhance diversity.
+High Quality: Pay attention to the quality of datasets (whether uni-modal answers are reasonable and accurate).
+Discriminative: Pay attention to the performance of this dataset on the model, and remove overly difficult subsets with little discrimination.
 3.2.2 Dataset Compression
-Regarding the existing large-scale uni-modal benchmarks, to reduce the evaluation cost of large-scale models, we designed a clustering-guided hierarchical sampling method.
+Regarding the existing large-scale uni-modal benchmarks, to reduce the evaluation cost of large-scale models, we designed a clustering-guided hierarchical sampling (CGHS) method as shown in Figure.2.
 This method achieves efficient compression and high evaluation consistency through the following steps:
 Question Characterization: Represent each question as an x-dimensional vector, where dimensions correspond to scores from different models on that question.
-Cluster-based Stratification: Utilize the K-Means algorithm to categorize questions into k clusters, each representing a "model performance similar" question type (e.g., easy questions, difficult questions, etc.).
+Cluster-based Stratification: Utilize the Kmeans++(Arthur and Vassilvitskii, 2007) algorithm to categorize questions into k clusters, each representing a "model performance similar" question type (e.g., easy questions, difficult questions, etc.).
 Hierarchical Sampling: Determine the sample size for each stratum based on cluster size proportions, and construct the final evaluation subset through simple random sampling.
@@
 Verify the consistency between model rankings and scores using Spearman Rank Correlation Coefficient (SRCC) and Pearson Linear Correlation Coefficient (PLCC).
-Finally, in order to ensure statistical stability, we repeat above steps by employing 5 random splits and 10-fold cross-validation.
-This method determines the optimal sample size through cost-benefit curve analysis, reducing evaluation costs by over 90% while maintaining accuracy (Figure.8).
+To ensure statistical stability, we repeat the above steps by using 5 random splits and performing 10-fold cross-validation.
+This approach identifies the optimal sample size via cost-benefit curve analysis, leading to a reduction in evaluation costs by over 90% while preserving accuracy, as shown in Figure.9.
 3.3 Multi-Step Open-Ended Questions
-Cross-modality problem requires the model to understand among multiple modalities with multi-step reasoning.
-However, existing evaluation benchmarks primarily focus on isolated problems where questions and answers remain independent of each other.
-Such a single-step evaluation paradigm cannot effectively assess the ability of a model to understand the complex, cross modality tasks that require multiple logical steps.
-Besides, most benchmarks adopt multiple-choice questions to verify the model’s performance, which reduces the difficulty of questions (e.g.
-allows for guessing) and may underestimate the model’s deficiencies in complex tasks.
-To address these issues, we introduce a new evaluation method, Multi-Step Open-Ended Questions (MO), which breaks down complex questions into multiple interdependent sub-tasks and requires the model to solve them step by step in the form of open-ended text.
-The construction process and example are shown in Figure.5.
-The multi-step open-ended question is worth 10 points.
-During the process of providing a reference answer, scores are assigned to each sub-question based on its difficulty and necessity for answering the final question, which facilitates the evaluation of the model’s accuracy at each step.
-Experiments (4.4.1) show that compared with single-step evaluation method (e.g.
+3.3.1 Question Type Definition
+Evaluating the multi-step reasoning capabilities of omni models presents a significant challenge.
+Real-world problems require models to integrate multi-modal information and execute a sequence of logical steps.
+However, current automated benchmarks, often relying on Outcome Reward Models (ORMs), typically provide only a binary pass/fail judgment.
+This approach fails to distinguish between a model that completes 80% of a task and one that fails at 20%, a crucial gap that human evaluators easily perceive.
+While alternatives like Process Reward Models (PRMs)Lightman et al.
+(2023) or multi-turn dialoguesReddy et al.
+(2019) exist, they are hampered by high implementation difficulty, low accuracy, or poor efficiency.
+Moreover, the prevalence of multiple-choice formats in existing benchmarks is unrepresentative of real-world, open-ended user queries and can mask model deficiencies.
+To address these issues, we propose an innovative Multi-Step Open-Ended Question (MO) type, designed for granular and realistic assessment.
+In the construction of MO dataset, complex problems are first deconstructed by human experts into a series of progressive, interdependent sub-questions.
+Each sub-question is assigned a score based on its importance, summing to a total of 10 points.
+During testing, all sub-questions are posed in a single turn, requiring the model to generate a step-by-step open-ended response.
+This method allows us to precisely quantify how far along a complex reasoning chain a model can proceed, offering a more accurate and insightful measure of its true capabilities.
+An example is shown in Figure.4.
+3.3.2 General Scoring Model
+Beside the dataset construction, multi-step open-ended question introduces a new challenge of automated evaluation.
+To overcome this obstacle, we propose a general scoring model that supports multi-choice question, single-step open-ended question and multi-step open-ended question at the same time.
+Since the task is to compare the target answer and the predicted answer, we use Qwen3-14B(Yang et al., 2025) as backbone and curate a training dataset as illustrated in Figure.5.
+One of the critical way to improve accuracy is to group questions into finer types and define appropriate criteria for each types.
+See details in Appendix.D.
+Through the human-in-the-loop dataset curation, the scoring model achieves 95% accuracy in out-of-distribution models and benchmarks.
+Experiments in Section.4.4.1 show that compared with single-step evaluation method (e.g.
 multiple-choice questions), multi-step open-ended questions can effectively observe the ability decay of models in long-chain reasoning, providing a more realistic difficulty for advanced models with stronger discrimination.
 Figure 4: Example of multi-step open-ended questions.
-Table 2: Cross-modal benchmark results on multimodal models
-Model Visual Audio Omni-MC Omni-MO
-MiniCPM-o-2.6 41.85 63.10 26.77 21.92
-Qwen-2.5-omni-3B 45.71 65.81 27.26 25.72
-Baichuan-omni-1.5 49.25 66.50 30.32 20.48
-Ming-lite-omni-1.5 48.68 71.70 32.10 26.08
-Qwen-2.5-omni-7B 52.81 68.11 33.23 29.12
-Gemini-2.0-Flash 66.80 74.41 46.26 40.96
-Gemini-2.5-Flash 76.64 75.20 68.23 62.00
-Gemini-2.5-Pro 80.10 78.70 76.61 68.84
-Table 3: Cross-modal analysis on model ability system
+Figure 5: Training dataset construction and model training pipeline for general scoring model.
+Table 2: Omni-modal benchmark results on multimodal models.
+Model Audio Visual Omni-MC Omni-MO
+Qwen-2.5-Omni-3B 54.40 42.67 27.80 24.76
+MiniCPM-O-2.6 56.50 42.27 28.60 23.76
+Ming-lite-Omni-1.5 58.30 46.28 28.90 25.48
+Baichuan-Omni-1.5 54.10 44.66 29.70 21.04
+Qwen-2.5-Omni-7B 60.20 50.68 32.60 27.72
+Qwen-3-Omni-30B-A3B 79.40 63.29 42.10 37.08
+Gemini-2.0-Flash 70.70 62.76 44.90 38.56
+Gemini-2.5-Flash 79.50 69.54 54.30 47.08
+Gemini-2.5-Pro 88.40 78.67 70.90 57.32
+Table 3: Omni-modal analysis on model ability taxonomy. Cross-modal Recognition refers to the set of other Perception capabilities except Cross-modal Alignment.
 Model Perception Reasoning Overall
@@
 Overall
-Sequential
+General
+Reasoning
+Temporal
 Reasoning
@@
 Reasoning
-General
-Reasoning
 Overall
-MiniCPM-o-2.6 25.55 36.11 27.75 38.78 22.94 26.36 26.40 26.77
-Qwen-2.5-omni-3B 27.74 50.00 32.37 38.78 27.52 22.09 25.28 27.26
-Baichuan-omni-1.5 27.74 47.22 31.79 28.57 35.78 27.52 29.75 30.32
-Ming-lite-omni-1.5 32.85 55.56 37.57 30.61 30.28 29.46 29.98 32.10
-Qwen-2.5-omni-7B 40.88 52.78 43.35 38.78 30.28 27.13 29.31 33.23
-Gemini-2.0-Flash 46.72 61.11 49.71 63.27 35.78 44.19 44.97 46.29
-Gemini-2.5-Flash 67.88 77.78 69.94 85.71 41.28 75.19 67.56 68.23
-Gemini-2.5-Pro 81.02 83.33 81.50 91.84 55.05 78.68 74.72 76.61
+Qwen-2.5-Omni-3B 29.84 35.94 33.09 20.65 50.00 20.83 23.98 27.80
+MiniCPM-O-2.6 26.70 30.88 28.92 26.62 42.42 26.67 28.40 28.60
+Ming-lite-Omni-1.5 28.80 35.94 32.60 24.38 43.94 24.17 26.53 28.90
+Baichuan-Omni-1.5 30.89 32.26 31.62 25.87 45.45 28.33 28.57 29.70
+Qwen-2.5-Omni-7B 38.22 36.41 37.25 28.11 43.94 26.67 29.59 32.60
+Qwen-3-Omni-30B-A3B 53.40 45.16 49.02 38.06 53.03 26.67 37.41 42.10
+Gemini-2.0-Flash 43.98 49.77 47.06 45.02 57.58 31.67 43.71 44.90
+Gemini-2.5-Flash 56.02 50.69 53.19 61.44 68.18 27.50 55.27 54.30
+Gemini-2.5-Pro 74.35 70.05 72.06 75.62 84.85 45.00 70.41 70.90
 4 Experiment and Analysis
 4.1 Experiment Setting
-We evaluate omni-models that support text, visual, and audio inputs simultaneously, including open-source models: Qwen-3-omni(Xu et al., 2025b), Qwen-2.5-omni-3B, Qwen-2.5-omni-7B(Xu et al., 2025a), Baichuan-omni-1.5(Li et al., 2025), MiniCPM-o-2.6(Yao et al., 2024), and Ming-lite-omni-1.5(AI et al., 2025), as well as closed-source models: Gemini-2.5-Pro, Gemini-2.5-Flash, and Gemini-2.0-Flash(Comanici et al., 2025).
-All model integrations strictly adhere to official implementations.
+We evaluate omni models that support text, visual, and audio inputs simultaneously, including open-source models: Qwen-3-Omni-30B-A3B-Instruct(Xu et al., 2025b), Qwen-2.5-Omni-3B, Qwen-2.5-Omni-7B(Xu et al., 2025a), Baichuan-Omni-1.5(Li et al., 2025), MiniCPM-O-2.6(Yao et al., 2024), and Ming-lite-Omni-1.5(AI et al., 2025), as well as closed-source models: Gemini-2.5-Pro, Gemini-2.5-Flash, and Gemini-2.0-Flash(Comanici et al., 2025).
+To have a fair comparison between instruct model and thinking model, we adopt similar way in Qwen-3(Xu et al., 2025b) that limits thinking budget to 128 tokens.
+We apply this restriction to Gemini-2.5-Pro and disable the thinking mode for both Gemini-2.5-Flash and Gemini-2.0-Flash.
+All the other model integrations strictly adhere to official implementations.
 In video processing, each model receives raw video and performs frame sampling according to its own sampling strategy.
-In the following sections, we conduct comprehensive experiments on MMAO-Bench and try to answer the following questions:
-1. How about the performance of current omni-models and their shortage abilities?
-2. What is the relationship of uni-modal ability and omni-modal ability?
-3. Could the MMAO-Bench provide an effective "All in One" evaluation for omni-models?
+In the following sections, we conduct comprehensive experiments on UNO-Bench and try to answer the following questions:
+  1. 1.
+    How about the performance of current omni models and their shortage abilities?
+  2. 2.
+    What is the relationship of uni-modal ability and omni-modal ability?
+  3. 3.
+    Could the UNO-Bench provide an effective evaluation for the omni model?
 4.2 Model Performance
 4.2.1 Overall Analysis
-As shown in Table 2, the proprietary Gemini series of models substantially outperform the evaluated open-source counterparts across all benchmarks.
-Gemini-2.5-Pro consistently leads, establishing a clear state-of-the-art.
-When analyzing the cross-modal tasks specifically, we observe distinct patterns.
-On the Omni-MC (Multiple-choice) benchmark, the performance of smaller open-source models like MiniCPM-o-2.6 (26.77) and Qwen-2.5-omni-3B (27.26) approximates that of random guessing (0.25).
-In contrast, larger models such as Qwen-2.5-omni-7B (33.23) demonstrate nascent capabilities beyond this baseline.
-Nevertheless, a significant performance gap remains, with the top open-source model trailing Gemini-2.5-Pro (76.61) by over 43 points.
-The Omni-MO (multi-step open-ended) benchmark, which demands more complex reasoning, proves to be even more challenging for all models.
-Performance universally drops compared to Omni-MC.
-Most open-source models score poorly, with the best among them, Qwen-2.5-omni-7B, achieving only 29.12.
-This indicates severe limitations in their ability to perform multi-step cross-modal reasoning.
-While the Gemini models also see a score reduction, they maintain a commanding lead, with Gemini-2.5-Pro (68.84) showcasing a far more robust, though not perfect, reasoning capability.
-This highlights that complex, multi-step, cross-modal understanding remains a critical and unresolved challenge, especially for the open-source community.
-In terms of perception capabilities, cross-modal alignment is more challenging than cross-modal recognition.
-Among open-source models, Qwen-2.5-omni-7B has the strongest alignment capability at 40.88, while Ming-lite-omni-1.5 has the strongest recognition capability at 55.56.
-The Gemini series significantly outperforms other models, and Gemini-2.5-Pro shows no significant weakness in alignment compared to its recognition capability, both exceeding 80.
-In terms of reasoning capabilities 3, spatial reasoning is currently the most challenging task, with the best-performing Gemini-2.5-Pro achieving only 55.05.
-Among open-source models, Baichuan-omni-1.5 performs the best with a score of 35.78.
-Ming-lite-omni-1.5 demonstrates the best performance in complex reasoning and general reasoning among open-source models.
-Overall, perception ability is relatively simpler compared to reasoning ability, and the model gap in reasoning ability is more significant.
-Taking Qwen-2.5-omni-7B and Gemini-2.5-Pro as examples, they differ by 38.2 points in perception and 45.41 points in reasoning.
-This demonstrates that reasoning ability is an extension of perception ability.
+Our main evaluation, summarized in Table.2, reveals a clear performance hierarchy where proprietary models, particularly Gemini-2.5-Pro, establish the state-of-the-art across all benchmarks.
+Meanwhile, progress within the open-source community is notable, with increased model scale and more training data, exemplified by Qwen-3-Omni-30B, leading to substantial improvements.
+Furthermore, we observe a strong positive correlation between a model’s performance on the foundational Audio and Visual tasks and its scores on the more demanding Omni benchmarks, suggesting that robust uni-modal perception is a prerequisite for advanced omni-modal understanding.
+On the Omni-MC (Multiple-choice) benchmark, which evaluates omni-modal comprehension, smaller open-source models exhibit performance marginally surpassing the random guess baseline (25.00), achieving scores between 27.80 and 29.70.
+The larger Qwen-3-Omni-30B marks a significant leap, with a score of 42.10 that approaches the performance of entry-level proprietary models like Gemini-2.0-Flash (44.90).
+Nevertheless, a substantial performance deficit persists when compared to the leading Gemini-2.5-Pro (70.90).
+This gap highlights the profound difficulty of advanced omni-modal comprehension, even in a multiple-choice format.
+The Omni-MO (Multi-Step Open-Ended) benchmark presents a considerably greater challenge, as evidenced by the universal and marked degradation in performance for all models relative to their Omni-MC scores.
+This degradation reveals a systemic limitation in multi-step omni-modal reasoning.
+For instance, the leading model, Gemini-2.5-Pro, attained a score of merely 57.32 on this benchmark, reflecting a decline of 13.58 points relative to its performance on the Omni-MC task.
+In comparison, the highest-scoring open-source model, Qwen-3-Omni-30B, achieved only 37.08 points.
+To dissect the models’ core capabilities, we perform a fine-grained analysis based on our proposed ability taxonomy, with detailed results presented in Table.3.
+In perception, a notable trend emerges: while smaller models find Recognition easier than Alignment, more powerful models like Qwen-3-Omni-30B-A3B and the Gemini-2.5 series exhibit stronger Alignment capabilities.
+This suggests that advanced models develop a more sophisticated grasp of inter-modal relationships.
+Among open-source models, Qwen-3-Omni-30B-A3B achieves the highest perception score (49.02).
+Gemini-2.5-Pro significantly leads overall, with both its Alignment (74.35) and Recognition (70.05) scores surpassing 70.
+In reasoning, Spatial Reasoning is consistently the most challenging task across all models.
+Even the top-performing Gemini-2.5-Pro only achieves 45.00.
+Notably, Baichuan-Omni-1.5 demonstrates the best spatial reasoning among open-source models with a score of 28.33.
+For General and Temporal Reasoning, the new Qwen-3-Omni-30B-A3B establishes itself as the open-source leader.
+Overall, reasoning proves to be a more challenging frontier than perception.
+This is highlighted by the performance gap between the leading proprietary model, Gemini-2.5-Pro, and the best open-source model, Qwen-3-Omni-30B-A3B.
+The disparity is 23.04 points in Perception (72.06 vs.
+49.02) but widens to a more substantial 33.00 points in Reasoning (70.41 vs.
+37.41).
+This indicates that advanced reasoning remains a key differentiator and a primary bottleneck for current multimodal models.
 4.2.2 Top-tier Analysis
@@
 On one hand, it stems from the leading uni-modal understanding ability.
-On the other hand, from the perspective of the reasoning process, Gemini possesses audio captioning capabilities, and it can also naturally incorporate audio content as a basis for reasoning.
+On the other hand, regarding the reasoning mechanism, Gemini is equipped with audio captioning functionalities as indicated in the technical report(Comanici et al., 2025), and illustrated in Figure.6.
+It can also naturally incorporate audio content as a foundation for reasoning.
 Existing open-source models, due to their smaller size, lack reasoning processes in a multimodal context.
 Limited reasoning mostly relies on text or images, with few involving specific audio content.
-However, does Gemini performance good enough? To answer this question, we invited human experts to conduct a competition.
+Figure 6: An example of omni-modal evaluation result. Gemini-2.5-Pro displays audio captions to assist in solving the problem.
+However, does Gemini performance good enough?
+To answer this question, we invited human experts to conduct a competition.
+It is important to note that, unlike the dataset annotators, the invited human experts have not been exposed to the question or the answer.
 Finding 1.
-Gemini-2.5-Pro has reached human comparable perception ability in omni-modal understanding, but still has a performance gap on reasoning ability.
+Gemini-2.5-Pro has reached human comparable perception ability in omni-modal understanding, yet there remains a gap in its reasoning performance.
 Compared to human experts, Gemini-2.5-Pro only has 8.3% less performance, which shows humanity comparable intelligence.
-The comparison of scores for specific ability items can be seen in Figure6.
-Dived into ability analysis, we find a interesting phenomena that human is more capable on reasoning than perception (81.3% > 74.3%), and perception ability is even lower than model’s.
-Through interviewing some of the human experts, we find that human may lost some information in video or audio in one hand, and in the other hand human has limited world knowledge than large language models.
+The comparison of scores for specific ability items can be seen in Figure.7.
+Upon examining ability analysis, we observe an intriguing phenomenon: humans are more proficient in reasoning as opposed to perception (81.3% compared to 74.3%), which contrasts with the model’s performance.
+By interviewing various human experts, it becomes evident that humans might miss some information in video or audio formats, and their world knowledge is more limited compared to large language models.
 Figure 7: The competition between human experts and Gemini-2.5-Pro. Gemini-2.5-Pro shows comparable perception capability but lower reasoning capability.
@@
 To investigate the relationship of uni-modal and omni-modal understanding ability, we conduct regression analysis and ablation experiments.
-Thanks to the unified ability systems and the high quality of omni-modal samples in MMAO-Bench, we find some interesting observations.
+Thanks to the unified ability taxonomy and the high quality of omni-modal samples in UNO-Bench, we find some interesting observations.
 Finding 2.
-Compositional Law: the performance of omni-modal ability is power-law related to the multiply of uni-modals’ performances.
-Observing the Table.2, we find that the omni-modal performances are correlation to the multiply of their uni-modals’ performances.
-In details, the performances of different models are relatively close in audio understanding but show significant differences in vision understanding, and these differences are also pronounced in the omni-modal considering both abilities.
-Gemini-2.5-Flash and Gemini-2.0-Flash are similar in audio understanding, but for vision modality, Gemini-2.5-Flash outperforms Gemini-2.0-Flash by 9.8, and it is 24.4 stronger in the cross-modal understanding.
-Finding 3.
-The weaker models show "short-board effect", while the stronger models demonstrate emergent ability.
-We draw the multiply of uni-modal scores as a dot line in the Figure.7.
-Compared with the cross-modal scores, models around the "random guessing stage" are below the line, while Gemini-2.5-Pro shows a significant gap above the line.
-Besides, these models demonstrate a power-law tendency that the stronger uni-modal ability can stimulate to much stronger cross-modal ability.
-This also emphasis that multiple modalities can enhance each other and lead to a higher intelligence.
+Compositional Law: the effectiveness of omni-modal capability is related to the product of the performances of individual modalities by a power-law.
+Observing the results in Table.2, we identify a strong correlation between a model’s omni-modal performance and its uni-modal capabilities.
+To formalize this, we derive a Compositional Law from a general functional form by applying two simplifying principles dictated by the omni-modal tasks proposed in our UNO-Bench.
+Let’s elaborate on the specifics below.
+General Model & Task Constraints.
+We begin by positing that the omni-modal performance $\mathcal{P}_{\text{Omni}}$ is a function of uni-modal performances $\mathcal{P}_{\text{A}}$ and $\mathcal{P}_{\text{V}}$.
+A general form can be written as:
+$\begin{split}\mathcal{P}_{\text{Omni}}(\mathcal{P}_{\text{A}},\mathcal{P}_{\text{V}})={}&f_{\text{A}}(\mathcal{P}_{\text{A}})+f_{\text{V}}(\mathcal{P}_{\text{V}})\\ &+f_{\text{I}}(\mathcal{P}_{\text{A}},\mathcal{P}_{\text{V}})+b\end{split}$ (1)
+where $f_{\text{A}},f_{\text{V}}$ represent modality independent path contributions, $f_{\text{I}}$ the interaction, and $b$ a baseline performance constant (e.g. random guess).
+We arrive at the following result through rigorous mathematical derivation, and the detailed derivation process is provided in the Appendix B.
+$\mathcal{P}_{\text{Omni}}=C\cdot\mathcal{P}_{\text{A}}^{\alpha}\mathcal{P}_{\text{V}}^{\beta}+b$ (2)
+where $C$ is a scaling constant, and exponents $\alpha,\beta$ model the interaction’s elasticity.
+We then posit a fusion symmetry assumption: in end-to-end omni models, the fusion mechanism does not inherently favor one modality over anotherXu et al.
+(2025b); Yao et al.
+(2024), implying symmetric scaling behavior.
+This leads to $\alpha=\beta$.
+Substituting this into Eq.
+10, we arrive at the Omni-modal Compositional Law:
+$\mathcal{P}_{\text{Omni}}=C\cdot(\mathcal{P}_{\text{A}}\times\mathcal{P}_{\text{V}})^{\alpha}+b$ (3)
+where $\alpha$ is the synergistic exponent, $C$ is a scaling coefficient, and $b$ is a baseline bias.
+A non-linear regression on data from leading models (Figure.8) yields the precise empirical formula:
+$\mathcal{P}_{\text{Omni}}\approx 1.0332\cdot(\mathcal{P}_{\text{A}}\times\mathcal{P}_{\text{V}})^{2.1918}+0.2422$ (4)
+This model demonstrates an exceptional fit, with a coefficient of determination ($R^{2}$) of 0.9759.
+Analysis of the fitted parameters reveals a clear transition from limited gains to emergent capabilities, driven by the super-linear nature of the law.
+Figure 8: The Compositional Law of Omni-modal Performance.
+Observed omni-modal scores (dots) versus the product of their uni-modal scores.
+The dashed line represents our fitted law (Eq.
+4), which shows a remarkable alignment with the empirical data ($R^{2}=0.9759$).
+The convex, accelerating curve visualizes the Power-law synergy.
+Power-law Synergy and Emergent Ability.
+The exponent $\alpha\approx 2.19$ is the most critical discovery, revealing a powerful Power-law synergy.
+Because $\alpha>1$, the function is convex, meaning the rate of performance gain accelerates.
+This explains the transition from a "short-board effect" to an "emergent ability" seen in Figure.8:
+  • Limited Gains at Low Performance: For models with weaker uni-modal abilities (e.g., MiniCPM-O), the curve is relatively flat.
+Small improvements in the product of uni-modal scores yield only marginal gains in omni-modal performance.
+This can be seen as a "short-board effect", where the system is not yet capable of effectively leveraging the combined inputs.
+  • Emergent Ability at High Performance: As uni-modal abilities strengthen (e.g., moving towards Gemini-2.5-Pro), the curve steepens dramatically.
+The same amount of improvement in the uni-modal product now yields a much larger increase in omni-modal performance.
+This accelerating return on investment is the quantitative signature of emergence, where stronger foundational skills unlock disproportionately powerful combined capabilities.
+Interpreting the Coefficients and Benchmark Coherence.
+The other parameters complete the picture.
+The bias term $b\approx 0.2422$ acts as a performance floor.
+As uni-modal performances approach zero, the system’s output converges to this value, which is strikingly close to the 0.25 random-guess accuracy of our benchmark.
+The scaling coefficient $C\approx 1.0332$, being remarkably close to unity, indicates a harmonious and naturally scaled system.
+We attribute this harmony not only to the models’ intrinsic fusion mechanisms but also to the coherent design of our benchmark itself.
+Additional fitted models are presented in Appendix.C.
+We argue that our proposed model is the most natural and interpretable among them.
+Interestingly, most models indicate that the visual modality tends to offer greater benefits than the audio modality.
+This phenomenon may be attributable to the relatively weaker visual capabilities of models at the current stage of development.
+It is worth emphasizing that this finding is directly attributed to the deliberate design of UNO-Bench.
+Specifically, it not only ensures a balanced distribution of capabilities across both uni-modal and omni-modal tasks, but also constructs the majority of questions to demand the joint processing of both modalities for resolution.
 Next, we conduct ablation studies to dig dive about the enhancement from vision and audio modality respectively.
-Figure 7: Relationship between uni-modal ability and cross-modal ability among omni-models.
 Table 4: Ablation of visual understanding ability.
 Model Perception Reasoning Overall
-Audio +Caption +Visual Audio +Caption +Visual Audio +Visual
-MiniCPM-o-2.6 26.01 30.06 27.75 24.83 25.28 26.40 25.16 26.77
-Qwen-2.5-omni-3B 21.39 29.48 32.37 22.60 24.83 25.28 22.26 27.26
-Baichuan-omni-1.5 22.54 30.06 31.79 28.86 31.54 29.75 27.10 30.32
-Ming-lite-omni-1.5 26.59 27.17 37.57 25.06 24.38 29.98 25.48 32.10
-Qwen-2.5-omni-7B 26.01 33.53 43.35 20.81 25.28 29.31 22.26 33.23
-Gemini-2.0-Flash 31.21 53.18 49.71 34.90 51.01 44.97 33.87 46.29
-Gemini-2.5-Flash 33.53 68.21 69.94 44.30 70.02 67.56 41.29 68.23
-Gemini-2.5-Pro 41.62 78.03 81.50 50.34 73.60 74.72 47.90 76.61
+Audio + Caption + Visual Audio + Caption + Visual Audio + Caption + Visual
+Qwen-2.5-Omni-3B 17.76 29.13 33.09 20.07 21.43 23.98 19.12 24.60 27.80
+MiniCPM-O-2.6 29.44 29.61 28.92 27.21 29.93 28.40 28.13 29.80 28.60
+Ming-lite-Omni-1.5 26.28 31.07 32.60 23.13 21.43 26.53 24.42 25.40 28.90
+Baichuan-Omni-1.5 22.14 32.04 31.62 23.81 26.70 28.57 23.12 28.90 29.70
+Qwen-2.5-Omni-7B 22.14 30.10 37.25 20.41 25.34 29.59 21.12 27.30 32.60
+Qwen-3-Omni-30B-A3B 27.01 46.84 49.02 18.71 39.63 37.41 22.12 42.60 42.10
+Gemini-2.0-Flash 25.55 44.17 47.06 29.76 45.58 43.71 28.03 45.00 44.90
+Gemini-2.5-Flash 22.63 49.03 53.19 29.08 53.23 55.27 26.43 51.50 54.30
+Gemini-2.5-Pro 37.71 63.83 72.06 42.18 65.99 70.41 40.34 65.10 70.90
 4.3.1 Ablation Visual Understanding
-To evaluate the enhancement of visual information on model understanding, we sequentially assessed the performance of Audio (only auditory information), +Caption (auditory + Caption information), and +Visual (auditory information + visual information), where captions were derived from descriptions of videos by Gemini-2.5-Pro.
-In the case of audio-only input, most models’ metrics are significantly lower than when both auditory and visual information are provided, and the metrics exhibit a random distribution around 25%.
-Gemini-2.5-Pro identified problem shortcuts through methods such as elimination, achieving a score of 47.9, which will be optimized in subsequent versions.
-After providing the caption information, all model metrics showed significant increases, but the magnitude of the rise varied considerably.
-In terms of perceptual ability, the Gemini series generally improved by 20 points, while the Qwen-2.5-omni series and Baichuan-omni-1.5 saw an improvement of over 8 points.
-The MiniCPM-o-2.6 showed a smaller increase of around 4 points, whereas the Ming-lite-omni-1.5 barely improved at all.
-This indicates that different models have varying abilities to reconstruct image content based on textual descriptions, and similar phenomena are observed in reasoning capabilities as well.
-Compared with caption information, the original visual input brings more information.
-On models with better baseline capabilities, such as Ming-lite-omni-1.5 and Qwen-2.5-omni-7B, caption information is far less than the original visual information.
-On models with slightly weaker baseline capabilities, the two are approximately the same.
-Interestingly, the Gemini series has a very high utilization rate of caption information, and even Gemini-2.0-Flash performs better with +caption than with +visual.
+To quantify the contribution of visual information, we conducted an ablation study with three settings: audio-only (Audio), audio plus high-quality textual captions of the visual scene (+ Caption), and the full audio-visual input (+ Visual).
+The captions were generated by Gemini-2.5-Pro to ensure descriptive richness.
+Results are detailed in Table.4.
+With only audio input, most models’ performance drops to a level near random guessing (around 20-28%), confirming the critical role of visual context.
+A notable exception is Gemini-2.5-Pro, which scores 40.34, suggesting an ability to leverage linguistic cues or shortcuts within the questions even without visual data.
+The introduction of Caption information yields significant but highly variable performance gains.
+Powerful models like the Gemini series and Qwen-3-Omni-30B-A3B demonstrate a substantial leap in performance (gains of 20-25 points), showcasing their strong ability to reconstruct scenes from textual descriptions.
+In contrast, models like MiniCPM-O-2.6 and Ming-lite-Omni-1.5 show minimal improvement, indicating a weaker capacity for this text-to-vision reasoning.
+Comparing Caption against full Visual input reveals a fascinating dichotomy.
+For the most capable model, Gemini-2.5-Pro, direct visual information provides a clear advantage over captions (70.90 vs.
+65.10), proving that raw visual data contains nuances that text cannot fully capture.
+However, for several other models, including Gemini-2.0-Flash and the powerful Qwen-3-Omni-30B-A3B, performance with captions is surprisingly on par with, or even slightly exceeds, that with direct visual input.
+This suggests that for these models, the language processing pathway may be more adept at extracting semantic meaning than their own visual encoders, highlighting a potential imbalance in their multimodal processing capabilities.
 4.3.2 Ablation Audio Understanding
-We categorize sound types into ambient sounds and dialogue, and provide their transcription versions based on Gemini-2.5-pro, where dialogue is given dialogue text, and ambient sounds are given sound descriptions.
-Similar to visual information, in the case of only providing auditory information, most models’ metrics are significantly lower than when both auditory and visual information are provided; after providing speech transcripts, the improvement is again more pronounced in the Gemini series.
-MiniCPM-o-2.6 performs better than +Audio on +ASR, confirming that the bottleneck of this model lies in audio understanding.
-Table 5: Ablation of audio understanding ability.
-Model Perception Reasoning Overall
-Video +ASR +Audio Video +ASR +Audio Video +ASR +Audio
-MiniCPM-o-2.6 27.75 31.79 27.75 27.07 29.98 26.40 27.26 30.48 26.77
-Qwen-2.5-omni-3B 27.17 30.06 32.37 18.79 23.94 25.28 21.13 25.65 27.26
-Baichuan-omni-1.5 26.01 27.17 31.79 27.07 23.04 29.75 26.77 24.19 30.32
-Ming-lite-omni-1.5 32.37 35.84 37.57 23.04 27.52 29.98 25.65 29.84 32.10
-Qwen-2.5-omni-7B 30.06 41.04 43.35 24.16 27.52 29.31 25.81 31.29 33.23
-Gemini-2.0-Flash 31.79 55.49 49.71 29.08 47.43 44.97 29.84 49.68 46.29
-Gemini-2.5-Flash 38.73 72.83 69.94 37.58 68.68 67.56 37.90 69.84 68.23
-Gemini-2.5-Pro 46.82 76.88 81.50 47.20 75.62 74.72 47.10 75.97 76.61
-In terms of environmental sound, most open-source models perform poorly.
-The two more capable models, Ming-lite-omni-1.5 and Qwen-2.5-omni-7B, show better performance in dialogue than in environmental sound.
-Additionally, we observed that the audio information of MiniCPM-o-2.6 and Ming-lite-omni-1.5 has side effects in environmental sound scenarios.
-When audio descriptions replace the original audio, both models show significant improvement, indicating that Ming-lite-omni-1.5 and Qwen-2.5-omni-7B have relatively weaker understanding capabilities of audio elements.
-In conversational scenarios, the Gemini series achieves ASR results that are nearly on par with, or even surpass, the original audio quality.
-This distinctive feature is absent in most other open-source models, with the exception of MiniCPM-o-2.6, where the +ASR results exceed +Audio 4 points.
+To isolate the impact of auditory information, we evaluated models under three conditions: visual-only (Visual), visual plus transcribed audio (+Caption), and the full audio-visual input (+Audio).
+We further divided the audio into two categories: Environment sounds, for which we provided textual descriptions, and Speech, for which we provided ASR transcripts.
+All transcriptions were generated by a powerful multimodal model.
+The results are presented in Table.5.
+The Visual-only setting results in significantly lower performance across all models, with Overall scores ranging from 21.20 to 33.70.
+This confirms the critical role of auditory context in multimodal understanding.
+The introduction of textual audio descriptions (+Caption) substantially boosts performance across the board.
+The improvement is particularly dramatic for high-capacity models like Gemini-2.5-Pro (+31.0 points Overall) and Qwen-3-Omni-30B-A3B (+17.4 points Overall), demonstrating their strong ability to integrate textual information.
+The comparison between +Caption and +Audio reveals crucial insights into the models’ raw audio processing capabilities.
+In environmental sound scenarios, understanding raw audio remains a significant challenge for most open-source models.
+For instance, Qwen-2.5-Omni-3B, MiniCPM-O-2.6, and Ming-lite-Omni-1.5 all exhibit considerably higher performance with textual descriptions (+Caption) than with the original audio (+Audio).
+This suggests that their audio encoders struggle to extract meaningful features from complex non-speech sounds, making them prefer clean textual summaries.
+In contrast, the most capable models—Gemini-2.5-Pro, Gemini-2.5-Flash, and Qwen-3-Omni-30B-A3B—demonstrate superior audio understanding by scoring higher in the +Audio setting, indicating they can extract richer information directly from the audio signal than is present in the provided caption.
+In conversational (Speech) scenarios, the results are more nuanced.
+The top-performing Gemini-2.5-Pro shows a substantial advantage with raw audio over ASR transcripts (+Audio 72.16 vs.
++Caption 66.00), indicating it effectively leverages paralinguistic cues such as tone, emotion, and prosody that are lost in transcription.
+Conversely, several other models, including the Qwen series and MiniCPM-O-2.6, perform slightly better with ASR transcripts (+Caption) than with raw audio.
+This points to a common bottleneck where imperfections in their audio encoders are a greater liability than the information lost during ASR, making clean text a more reliable input.
+Notably, Gemini-2.5-Flash achieves nearly identical scores in both settings, suggesting its ASR and audio understanding capabilities are exceptionally well-aligned.
 Table 5: Ablation of audio understanding ability.
 Model Environment Speech Overall
-Visual +Caption +Audio Visual +ASR +Audio Visual +ASR/Caption +Audio
-MiniCPM-o-2.6 33.33 31.11 28.89 26.70 30.54 26.53 27.26 30.48 26.77
-Qwen-2.5-omni-3B 15.56 31.11 28.89 21.64 25.31 27.23 21.13 25.65 27.26
-Baichuan-omni-1.5 26.67 22.22 33.33 26.70 24.43 30.02 26.77 24.19 30.32
-Ming-lite-omni-1.5 35.56 37.78 28.89 24.96 29.32 32.29 25.65 29.84 32.10
-Qwen-2.5-omni-7B 28.89 40.00 26.67 25.65 30.72 33.86 25.81 31.29 33.23
-Gemini-2.0-Flash 37.78 42.22 40.00 29.32 50.44 46.77 29.84 49.68 46.29
-Gemini-2.5-Flash 53.33 62.22 51.11 36.47 70.51 69.63 37.90 69.84 68.23
-Gemini-2.5-Pro 57.78 71.11 75.56 46.07 76.44 76.79 47.10 75.97 76.61
+Visual +Caption +Audio Visual +Caption +Audio Visual +Caption +Audio
+Qwen-2.5-Omni-3B 26.28 41.03 34.62 24.76 26.66 26.54 25.00 28.90 27.80
+MiniCPM-O-2.6 26.92 39.74 34.62 28.08 28.44 27.49 27.90 30.20 28.60
+Ming-lite-Omni-1.5 31.41 43.59 35.26 22.27 25.59 27.73 23.70 28.40 28.90
+Baichuan-Omni-1.5 25.64 32.05 28.85 23.70 23.58 29.86 24.00 24.90 29.70
+Qwen-2.5-Omni-7B 30.77 41.03 37.18 24.41 33.06 31.75 25.40 34.30 32.60
+Qwen-3-Omni-30B-A3B 32.05 48.08 48.72 23.58 41.23 40.88 24.90 42.30 42.10
+Gemini-2.0-Flash 25.00 48.08 45.51 22.87 48.93 44.79 23.20 48.80 44.90
+Gemini-2.5-Flash 17.95 48.72 49.36 21.80 55.09 55.21 21.20 54.10 54.30
+Gemini-2.5-Pro 32.69 57.69 64.10 33.89 66.00 72.16 33.70 64.70 70.90
 4.4 Benchmark Analysis
-In this section, we verify the effectiveness of MMAO-Bench on three aspects, the performance of multi-step open-ended question, the performance of dataset compression and the benchmark comparison with other open-sourced benchmarks.
+In this section, we verify the effectiveness of UNO-Bench on three aspects, the performance of multi-step open-ended question, the performance of dataset compression and the benchmark comparison with other open-source benchmarks.
 4.4.1 Multi-Step Open-Ended Question Analysis
 In this work, we introduce a new type of evaluation method, multi-step open-ended question, which effectively assess the complex reasoning ability, especially appears in cross-modality understanding.
-As shown in Table.7, the experimental results demonstrate the stratified performance among different models.
-Gemini-2.5-Pro leads with a comprehensive score of 68.84, showcasing its outstanding capabilities as a top-tier closed-source model.
-Following closely are Gemini-2.5-Flash and Gemini-2.0-Flash, which also form a clear leading tier.
-Among open-source models, Qwen-2.5-omni-7B performs best, significantly outperforming other smaller-scale open-source models.
-This indicates that model scale and advanced architecture remain key factors in addressing complex multi-question tasks.
-As the depth of questions increases, model performance generally declines.
-This is a very significant trend, also validating the effectiveness of our dataset design.
-For almost all models, their performance on small question 1 is better than on small question 2, while the score for small question 3+ is the lowest.
-For example, Gemini-2.5-Pro’s score on the "perception" task dropped from 26.60 on small question 1 to 21.79 on small question 3+; on the "reasoning" task, the score sharply decreased from 38.2 on small question 1 to 21.33 on small question 3+.
-This performance decay phenomenon reveals the common challenges current large models face in handling long-range dependencies, maintaining conversational context, and conducting multi-step chain-of-thought reasoning.
-Initial perception and simple reasoning are relatively easy, but as conversations deepen, the accumulation of errors and loss of context lead to significantly lower success rates in answering more complex subsequent questions.
-Reasoning ability is the key bottleneck that distinguishes model performance.
-Comparing the total scores of "perception" and "reasoning" dimensions, we found that all models scored higher on perception tasks than on reasoning tasks.
-For example, Gemini-2.5-Pro’s perception total score is 78.4, while its reasoning total score is 66.45.
-Among open-source models, Qwen-2.5-omni-7B’s perception total score (40.8) is significantly higher than its reasoning total score (26.20).
-This indicates that although models have already acquired basic perception capabilities (such as object recognition and sound recognition), there is still significant room for improvement in their ability to transform this perceptual information into complex logical, spatial, or causal reasoning.
-Our dataset successfully amplifies this capability gap through its progressive question design.
+As shown in Table.6, the experimental results on our multi-step open-ended questions reveal a clear performance stratification among models.
+Gemini-2.5-Pro establishes itself as the top-tier model with an overall score of 57.32, with Gemini-2.5-Flash (47.08) and Gemini-2.0-Flash (38.56) forming a distinct second tier.
+Among open-source models, Qwen-3-Omni-30B-A3B emerges as the clear leader with a score of 37.08, significantly outperforming smaller-scale models like Qwen-2.5-Omni-7B (27.72).
+This starkly illustrates that both advanced architecture and model scale are pivotal factors for success in complex, multi-turn multimodal tasks.
+As the depth of questions increases from Q1 to Q3+, most models exhibit a general decline in performance, confirming the effectiveness of our dataset’s progressive difficulty.
+For instance, the leading open-source model, Qwen-3-Omni-30B-A3B, sees its overall score drop from 18.08 on the first question (Q1) to 14.18 (Q2) and further to 11.42 (Q3+).
+This decay highlights a common challenge for current models in handling long-range dependencies, maintaining conversational context, and performing multi-step reasoning.
+However, a notable exception is Gemini-2.5-Pro, whose performance on the second question (Q2) surpasses its score on the first (24.48 vs.
+23.44), before declining on subsequent questions.
+This unique pattern suggests a superior ability to utilize the context from the initial turn to enhance its understanding and response in the subsequent turn, a capability not observed in other models.
+Reasoning ability remains the key bottleneck that differentiates model performance.
+For all open-source models and the lower-tier Gemini models, scores on Perception tasks are considerably higher than on Reasoning tasks.
+The gap is particularly pronounced for Qwen-3-Omni-30B-A3B, which scores 53.8 in Perception but only 32.9 in Reasoning.
+This indicates that while these models have developed solid foundational perception capabilities, converting this perceptual input into complex logical or causal reasoning remains a major hurdle.
+Interestingly, Gemini-2.5-Pro is the only model that defies this trend, achieving a higher score in Reasoning (58.1) than in Perception (54.2).
+This exceptional result demonstrates that state-of-the-art models are beginning to overcome the reasoning bottleneck, showcasing advanced cognitive abilities that are on par with, or even exceed, their perceptual skills.
+The design of our dataset successfully magnifies this critical capability gap between the SOTA and other models.
 Table 6: Performance on Multi-Step Open-Ended Questions.
@@
 Q1 Q2 Q3+ All Q1 Q2 Q3+ All Q1 Q2 Q3+ All
-Baichuan-omni-3B 8.95 7.75 5.38 24.2 11.8 9.6 9.33 19.55 9.52 8.12 5.87 20.48
-MiniCPM-6-2.6B 10.65 9.35 8.11 34.2 22.4 11.8 6.67 18.85 11.16 8.12 6.33 21.92
-Qwen-2.5-omni-3B 10.95 8.59 8.76 34.2 19.4 13.2 5.33 23.60 12.64 9.52 8.33 25.72
-Ming-lite-omni 8.35 7.20 6.29 33.2 20.8 12.4 4.67 24.30 12.68 9.96 7.69 26.08
-Qwen-2.5-omni-7B 12.85 9.75 7.90 40.8 18.4 17.6 16.00 26.20 13.96 11.33 8.92 29.12
-Gemini-2.0-Flash 17.55 13.60 11.32 56.2 27.8 22.0 21.33 37.15 19.6 15.28 12.56 40.96
-Gemini-2.5-Flash 23.40 25.05 20.66 72.4 35.2 31.4 26.00 59.40 25.76 26.32 21.32 62.00
-Gemini-2.5-Pro 26.60 29.10 21.79 78.4 38.2 33.8 21.33 66.45 28.92 30.04 21.74 68.84
+Baichuan-Omni-1.5 15.4 8.2 5.33 25.2 9 7.25 5.75 18.9 10.28 7.44 5.7 20.16
+MiniCPM-O-2.6 20.0 6.2 11.33 29.6 9.05 9.55 8.02 22.3 11.24 8.88 8.43 23.76
+Qwen-2.5-Omni-3B 19.8 12.2 5.33 33.6 10.7 7.2 8.86 22.55 12.52 8.2 8.42 24.76
+Ming-lite-Omni-1.5 19.6 12.4 4.67 33.4 10.9 8.4 7.92 23.5 12.64 9.2 7.52 25.48
+Qwen-2.5-Omni-7B 20.2 15.0 12.0 38.8 12.15 8.99 7.83 24.95 13.76 10.2 8.35 27.72
+Qwen-3-Omni-30B-A3B 25.0 22.8 20.0 53.8 16.35 12.01 10.19 32.9 18.08 14.18 11.42 37.08
+Gemini-2.0-Flash 25.2 19.4 14.67 49.0 15.5 14.05 13.02 35.95 17.44 15.12 13.22 38.56
+Gemini-2.5-Flash 31.6 22.6 12.0 57.8 18.35 17.35 16.42 44.4 21.0 18.4 15.87 47.08
+Gemini-2.5-Pro 25.6 22.2 21.33 54.2 22.9 25.05 19.43 58.1 23.44 24.48 19.67 57.32
 4.4.2 Dataset Compression
@@
 To evaluate the consistency of model ranking and the best size of compression data size, we conduct several experiments to analysis.
-Baseline dataset consists of  8000 samples including 15 open-source benchmarks (e.g.
-MathVista and MMAU, details see appendix) and 12 models evaluation results on them, which splits into 8/4 on models as training/test set.
-Kmeans is used to cluster with K=48.
+The baseline data set consists of  8000 samples including 18 open-source benchmarks(e.g.
+MathVista and MMAU, details see Appendix.A) and 20 models evaluation results on them, which split into 12/8 on models as training/test set.
+Kmeans++(Arthur and Vassilvitskii, 2007) is used to cluster with K=48.
 To eliminate the random factor, we conduct 5-fold settings and evaluate 10 times on each setting.
-To verify the compression performance, we use metrics Margin of Error(MoE), Root Mean Square Error(RMSE), CI Coverage, Spearman’s linear correlation coefficient(SRCC) and Pearson’s linear correlation coefficient(PLCC).
-The experimental result is shown in Figure.8, when compression rate is 10%, the coverage is nearly 100% and PLCC is above 0.98, SRCC is above 0.98, RMSE is 0.06, which means the model ranking is consistent.
+To verify the compression performance, we define these metrics: Spearman’s Rank Correlation Coefficient(SRCC) for ranking consistency, Pearson’s Linear Correlation Coefficient(PLCC) for linear value consistency, Root Mean Square Error(RMSE) for numerical precision, Margin of Error(MoE) for quantifying estimation uncertainty, and Confidence Interval Coverage(CIC) for statistical reliability.
+The experimental result is shown in Figure.9.
+At a 10% sampling rate, our method achieved excellent results on test-set.
+Both SRCC and PLCC exceeded 0.98, indicating near-perfect preservation of ranking and value relationships.
+The RMSE was below 0.02 with a corresponding MoE of 0.024; together, these values signify high numerical precision and a tight estimation range.
+Furthermore, the CIC was approximately 95%, confirming the statistical unbiasedness of the sample.
 Figure 9: Data compression performance.
 4.4.3 Benchmark Comparison
-Besides dataset quality and evaluation efficiency, a good benchmark should provide sufficient metric discriminability and a progressive difficulty scale to drive model capability growth.
-Compared with existing omni-modal benchmarks (including OmniBench(Li et al., 2024b), WorldSense(Hong et al., 2025) and AVOdyssey(Gong et al., 2024)), MMAO-Bench has relatively consistent rankings but a higher standard deviation of 19.5 in evaluation scores, while the best other OmniBench only has 12.75, showing the MMAO-Bench has greater discriminative ability.
-A simple way to improve discrimination is increasing problem difficulty, which can’t guide model’s growth.
-As a verification, we can see that Gemini-2.5-pro scored 76.61 on MMAO-Bench, but only 65.49 on OmniBench, indicating that MMAO-Bench has appropriate difficulty levels.
+To ensure the quality of dataset, we conduct quality check on 10%-20% random samples in each benchmarks.
+As shown in Table.1, UNO-Bench has 100% accuracy on omni-modal dataset while 98% questions requires cross-modality to solve.
+It shows the highest quality among existing omni model benchmarks.
+An effective benchmark must provide both a clear performance ladder and a meaningful difficulty range, and our UNO-Bench is engineered to deliver on both fronts as shown in Figure.10.
+It excels in discriminability, establishing substantial and remarkably linear intervals of ~10-12 points between adjacent models.
+This superior discriminability comes from a well-calibrated difficulty.
+UNO-Bench creates a vast 31.9 point performance gap between the top and bottom models, effectively separating their capabilities.
+This approach avoids the pitfall of being universally difficult, a problem seen in AV-Odyssey where all models are compressed into a narrow, low-scoring band (34-45).
+By combining a structured performance ladder with a balanced challenge, UNO-Bench serves as a more reliable and insightful tool for gauging genuine progress in the field.
 Figure 10: Comparison between omni benchmarks.
@@
 1.
-With unified data framework in MMAO-Bench, we found that the multimodal capability may not simply be a linear superposition of uni-modalities, but rather follows a significant multiplicative relationship.
+With unified data framework in UNO-Bench, we found that the multimodal capability may not simply be a linear superposition of uni-modalities, but rather follows a significant multiplicative relationship.
 The evaluation results show that it manifests as a bottleneck effect on weak models, while exhibiting synergistic promotion on strong models.
 2.
-Both uni-modal and omni-modal understanding capability of the Gemini series far surpasses existing open-source omni-models.
+Both uni-modal and omni-modal understanding capability of the Gemini series far surpasses existing open-source omni models.
 The Gemini-2.5-Pro shows comparable perception capability with human experts but still has a performance gap in reasoning aspect.
-3. Besides better dataset quality and evaluation efficiency, MMAO-Bench can provide sufficient metric discriminability and a progressive difficulty scale to drive model capability growth.
-In the future work, we will extend the dataset’s scale by human-in-the-loop automatic pipeline and hold a private test set to avoid hacking.
+3. Besides better dataset quality and evaluation efficiency, UNO-Bench can provide sufficient metric discriminability and a progressive difficulty scale to drive model capability growth.
+In the future work, we will extend the dataset’s scale by human-in-the-loop automated pipeline and hold a private test set to avoid hacking.
 The ability coverage also needs to extend to more difficult reasoning tasks like STEM and code.
 At the same time, the relationship among cross-modals understandings and how to improve them are still exciting problem to explore.
+Furthermore, our compositional law has only been validated on UNO-Bench with uniformly distributed capabilities.
+Whether this law still holds under different capability distributions remains to be explored.
+Acknowledgement
+We hereby express our appreciation to the LongCat Team EVA Committee for their valuable assistance, guidance, and suggestions throughout the course of this work.
+Appendix A Benchmarks Utilized in Dataset Compression
+To construct our compressed datasets, we utilized a variety of benchmarks for both visual and audio modalities.
+For the visual component, we curated data from 15 public and several in-house benchmarks that assess a range of capabilities, including general visual question answering, document and chart comprehension, STEM/scientific reasoning, and multi-image understanding.
+For the audio component, we used 3 audio question answering benchmarks.
+The detailed composition of the resulting uni-modal dataset is presented in Figure.11.
+Figure 11: The distribution of the uni-modal benchmarks in UNO-Bench.
+In addition to publicly available benchmarks, we incorporated several in-house benchmarks both before and after compression to ensure the reasonableness of the data distribution.
+  • General visual question answering, RealWorldQAxAI (2023), MMEChaoyou et al. (2023), SeedBenchLi et al. (2023).
+  • Document and chart understanding, OCRBench Liu et al. (2024d), FoxLiu et al. (2024a), DocLocal4kYe et al. (2023).
+  • Stem & reasoning, MMMUYue et al.
+(2024), MMMU-ProYue et al.
+(2025), CMMMUGe et al.
+(2024), MathVistaLu et al.
+(2024b), MathVisionWang et al.
+(2024b), ScienceVistaTeam et al.
+(2025), GMAI-MMBenchYe et al.
+(2024).
+  • Multi-image Understanding, ReMiKazemi et al. (2024), MuirBenchWang et al. (2024a).
+  • Audio question answering, MMAUSakshi et al. (2025), MMSUWang et al. (2025), SDQAFaisal et al. (2021).
+Appendix B Rigorous Derivation of the Compositional Law
+Defining the performance gain as $\mathcal{P}^{\prime}_{\text{Omni}}=\mathcal{P}_{\text{Omni}}-b$. From Eq. 1, we have:
+$\mathcal{P}^{\prime}_{\text{Omni}}(\mathcal{P}_{\text{A}},\mathcal{P}_{\text{V}})=f_{\text{A}}(\mathcal{P}_{\text{A}})+f_{\text{V}}(\mathcal{P}_{\text{V}})+f_{\text{I}}(\mathcal{P}_{\text{A}},\mathcal{P}_{\text{V}})$ (5)
+Due to the high quality of our benchmark, where a task is unsolvable if either modality is absent, causing the performance to drop to its baseline (e.g.
+random guessing).
+we can have a strict boundary condition:
+$\mathcal{P^{\prime}}_{\text{Omni}}(\mathcal{P}_{\text{A}},0)=0$ and (6)
+$\mathcal{P^{\prime}}_{\text{Omni}}(0,\mathcal{P}_{\text{V}})=0$ and
+$\mathcal{P^{\prime}}_{\text{Omni}}(0,0)=0$
+Applying the boundary condition of Eq. 6 to Eq. 5, we find that the gain is a second-order mixed difference of $f_{\text{I}}$:
+$\begin{split}\mathcal{P}^{\prime}_{\text{Omni}}(\mathcal{P}_{\text{A}},\mathcal{P}_{\text{V}})=&f_{\text{I}}(\mathcal{P}_{\text{A}},\mathcal{P}_{\text{V}})-f_{\text{I}}(\mathcal{P}_{\text{A}},0)\\ &-f_{\text{I}}(0,\mathcal{P}_{\text{V}})+f_{\text{I}}(0,0)\end{split}$ (7)
+Substituting the Taylor series of $f_{\text{I}}$ around $(0,0)$ into Eq. 7, the performance gain is thus exactly equal to the sum of all pure interaction terms from $f_{\text{I}}$:
+$\begin{split}\mathcal{P}^{\prime}_{\text{Omni}}(\mathcal{P}_{\text{A}},\mathcal{P}_{\text{V}})&=\sum_{i\geq 1,j\geq 1}c_{ij}\mathcal{P}_{\text{A}}^{i}\mathcal{P}_{\text{V}}^{j}\end{split}$ (8)
+where the coefficients $c_{ij}$ are constants derived from the partial derivatives of $f_{\text{I}}$ at the origin.
+For sufficiently small uni-modal performances, we can approximate this series by its leading-order term:
+$\mathcal{P}^{\prime}_{\text{Omni}}(\mathcal{P}_{\text{A}},\mathcal{P}_{\text{V}})\approx c_{11}\mathcal{P}_{\text{A}}\mathcal{P}_{\text{V}}$ (9)
+This result strongly motivates modeling the interaction with the general multiplicative Cobb-Douglas form. Re-introducing the baseline $b$ yields our final Compositional Law:
+$\mathcal{P}_{\text{Omni}}=C\cdot\mathcal{P}_{\text{A}}^{\alpha}\mathcal{P}_{\text{V}}^{\beta}+b$ (10)
+Appendix C Model Selection for the Compositional Law
+To validate our choice of the Compositional Law, we compared its performance against several alternative models.
+The fitting results for all candidate models on our 9-model dataset are summarized in Table.7.
+Table 7: Fitting results for all candidate models.
+While more complex models achieve higher fitting scores ($R^{2}$), their parameters lack physical interpretability (e.g., negative exponents), indicating severe overfitting on our small dataset.
+Our chosen Symmetric Power Law offers the best balance of a high $R^{2}$ value and theoretical soundness.
+Model Name $R^{2}$ RMSE Fitted Equation
+Generalized Power Law 0.999 0.005 $P_{Omni}\approx 1.33\cdot P_{Audio}^{-1.59}\cdot P_{Visual}^{5.09}+0.24$
+Linear Interaction 0.995 0.010 $P_{Omni}\approx 0.97-2.01P_{Audio}-0.59P_{Visual}+2.85(P_{Audio}\times P_{Visual})$
+Weighted Sum Power Law 0.995 0.010 $P_{Omni}\approx 1.19\cdot(-0.20P_{Audio}+1.20P_{Visual})^{3.83}+0.24$
+Symmetric Power Law 0.976 0.022 $\boldsymbol{P_{Omni}\approx 1.03\cdot(P_{Audio}\times P_{Visual})^{2.19}+0.24}$
+Simple Linear 0.945 0.033 $P_{Omni}\approx-0.15-0.37P_{Audio}+1.43P_{Visual}$
+As shown in Table.7, more complex models like the ‘Generalized Power Law’ achieve a near-perfect fit on the training data.
+However, this superior performance is misleading.
+These models yield parameters that are physically implausible, such as negative exponents (e.g., $P_{Audio}^{-1.59}$) or negative weights.
+Such parameters would illogically imply that improving a model’s uni-modal capability could degrade its omni-modal performance.
+This is a classic symptom of overfitting, where a model with high capacity learns the noise in a small dataset rather than a generalizable underlying trend.
+In contrast, our proposed Symmetric Power Law provides an excellent fit ($R^{2}=0.976$) while maintaining theoretical coherence.
+All its parameters are positive and have clear interpretations: a super-linear synergy ($\alpha=2.19>1$) between modalities, a positive scaling factor ($C=1.03$), and a reasonable baseline score ($b=0.24$).
+Following the principle of Occam’s Razor, we select this model as it offers the most parsimonious, robust, and interpretable explanation for the observed phenomenon.
+Interestingly, while the parameters from the overfitted models are invalid, they consistently suggest a stronger influence from the visual modality (e.g., the large positive exponent for $P_{Visual}$ in the ‘Generalized Power Law‘).
+This hints that while our symmetric law captures the primary collaborative effect, the visual component may play a slightly more dominant role, a direction for future investigation.
+Appendix D General Scoring Model
+One of the critical way to improve accuracy is to group questions into finer types and define appropriate criteria for each types. We define six different question types in Figure.12.
+Figure 12: Definition of finer question types for general scoring model.
+Appendix E Model Ability Taxonomy
+This section will provide specific definitions for each ability item and present examples of various task types.
+The specific model abilities and task types of the Perception dimension can be seen in Figure.13, and the Reasoning dimension can be seen in Figure.14.
+Figure 13: Definition of the Perception Dimension.
+Figure 14: Definition of the Reasoning Dimension.
+Specific examples are provided for each model ability.
+Examples of Perception ability including Object Perception, Spatial Perception, Cross-Modal Alignment, Attribute Perception, Scenario Perception, Cross-Modal Conversion and Semantic Understanding can be seen in Figure.15.
+Examples of Reasoning ability including Complex Reasoning, Temporal Reasoning, Spatial Reasoning, Life Reasoning, STEM Reasoning and Code can be respectively seen in Figure.16.
+Figure 15: Example of each ability in perception dimension.
+Figure 16: Example of each ability in reason dimension.
+Appendix F Omni-modal Solubility and Uni-modal Solubility Cases
+Omni-modal solubility means that both visual and audio information are required to answer the question.
+While uni-modal solubility means that only visual or audio information is required to answer the question.
+There are two uni-modal solubility cases.
+One is when audio and video present the same amount of information, and only one modality is needed to solve the problem, as shown in Figure.17(b).
+The other is when one modality provides useful information, while the other modality is unhelpful, as shown in Figure.17(c).
+The omni-modal solubility case can be seen in Figure.17(a).
+Figure 17: Comparison of the problem solvable by each modality.

### v3

@@
 However, the correlation between uni-modal and omni-modal remains unclear, which requires comprehensive evaluation to drive omni model’s intelligence evolution.
-In this work, we propose a novel, high quality and UNified Omni model benchmark, UNO-Bench, which effectively assesses both UNi-modal and Omni-modal capabilities.
-The benchmark consists of 3730 human curated samples, with 98% cross-modality solvability, across 44 task types, and an innovative multi-step open-ended question type for assessing complex reasoning.
-Besides, a general scoring model supporting 6 question types is proposed for automated evaluation with 95% accuracy.
+In this work, we introduce a novel, high-quality, and UNified Omni model benchmark, UNO-Bench.
+This benchmark is designed to effectively evaluate both UNi-modal and Omni-modal capabilities under a unified ability taxonomy, spanning 44 task types and 5 modality combinations.
+It includes 1250 human curated samples for omni-modal with 98% cross-modality solvability, and 2480 enhanced uni-modal samples.
+The human-generated dataset is well-suited to real-world scenarios, particularly within the Chinese context, whereas the automatically compressed dataset offers a 90% increase in speed and maintains 98% consistency across 18 public benchmarks.
+In addition to traditional multi-choice questions, we propose an innovative multi-step open-ended question format to assess complex reasoning.
+A general scoring model is incorporated, supporting 6 question types for automated evaluation with 95% accuracy.
 Experimental result shows the Compositional Law between omni-modal and uni-modal performance and the omni-modal capability manifests as a bottleneck effect on weak models, while exhibiting synergistic promotion on strong models.
-Our code and data are available at GitHub.
-UNO-Bench: A Unified Benchmark for Exploring the Compositional Law Between Uni-modal and Omni-modal in OmniModels
-Chen Chenthanks: Equal contribution.
-Zeyang Hu11footnotemark: 1   Fengjiao Chenthanks: Project Lead.
-Liya Ma   Jiaxing Liu   Xiaoyu Li22footnotemark: 2   Xuezhi Cao Meituan, Beijing, China.
-{chenchen165, huzeyang, chenfengjiao02, maliya06}@meituan.com {liujiaxing10, lixiaoyu28, caoxuezhi}@meituan.com
+GitHub: https://github.com/meituan-longcat/UNO-Bench
+Hugging Face: https://huggingface.co/datasets/meituan-longcat/UNO-Bench
+Figure 1: Benchmark Statistics and Evaluation Results.
 1 Introduction
@@
 Existing omni model evaluation benchmarks remain relatively scarce and focus on different aspects.
-For example, some prioritize image comprehension(Li et al., 2024b), others emphasize video understanding(Hong et al., 2025), while a subset concentrates on speech interaction(Gong et al., 2024).
+For example, some prioritize image comprehension(Li et al., 2024a), others emphasize video understanding(Hong et al., 2025), while a subset concentrates on speech interaction(Gong et al., 2024).
 Notably, existing datasets are exclusively English-centric, lacking evaluation benchmarks for Chinese linguistic contexts.
-The ideal omni model should simultaneously preserve visual understanding capabilities (e.g., MMBench(Liu et al., 2024b)/MathVista(Lu et al., 2024a)/MVBench(Li et al., 2024a)), speech interaction proficiency (e.g., MMAU(Sakshi et al., 2025)), and cross-modal integration capacity (e.g., OmniBench(Li et al., 2024b)/WorldSense(Hong et al., 2025)).
+The ideal omni model should simultaneously preserve visual understanding capabilities (e.g., MMBench(Liu et al., 2024a)/MathVista(Lu et al., 2024a)/MVBench(Li et al., 2024b)), speech interaction proficiency (e.g., MMAU(Sakshi et al., 2025)), and cross-modal integration capacity (e.g., OmniBench(Li et al., 2024a)/WorldSense(Hong et al., 2025)).
 Current evaluation paradigms employ disjointed benchmark suites for separate capability testing, creating resource-intensive evaluation processes and disconnected modality assessments.
@@
 These issues limit the evaluation and analysis of omni models’ capabilities.
-Dataset Omni-modal Uni-modal Acc. Solvable Source #Tasks #QA Pairs QA Type Language
-MMBench I - - 80% private 20 3217 MC EN/CH
-MMAU A - - 15% private 27 10000 MC EN
-MVBench V - - public 20 4000 MC EN
-OmniBench I+A 75% 90% public 8 1142 MC EN
-AV-Odyssey I+V+A 91% 99% public 26 4555 MC EN
-WorldSense V+A 99% 23% public 26 3172 MC EN
-Daily-Omni V+A 94% 59% public 6 1197 MC EN
-UNO-Bench-omni I+V+A -  100% 98% 90% private 44 1250 MC/MO EN/CH
-UNO-Bench-uni - I/V/A 86% 97% 42% private 44 2480 MC EN/CH
-Table 1: Comparison of MultiModal Benchmarks, with I, A, V, and T representing image, audio, video, and text modalities, respectively.
-It reports on the accuracy of question-answer pairs and the percentage of questions requiring omni-modal solutions, labeled as Acc.
-and Solvable.
-The Source category specifies the origin of the materials.
-Private sources, as opposed to public ones, can prevent data contamination.
-QA types include MC for multi-choice questions and MO for multi-step open-ended questions.
-EN and CH denote English and Chinese languages.
-UNO-Bench includes 1250 human-curated samples in the omni-modal section (referred to as -omni) and 2480 enhanced samples in the uni-modal section (referred to as -uni).
 Due to the limited quality and coverage of existing benchmarks, we introduce a novel and unified benchmark UNO-Bench.
@@
 Establish a high quality and diversity dataset construction pipeline including human-centric process and automated data compression.
-As as result, UNO-Bench comprises 1250 human curated samples for omni-modal with 98% cross-modality solvability, and 2480 enhanced samples for uni-modal, across 44 task types and 6 question types.
+As a result, UNO-Bench comprises 1250 human curated samples for omni-modal with 98% cross-modality solvability, and 2480 enhanced samples for uni-modal, across 44 task types and 5 modality combinations.
 The human-created novel dataset is well-suited to real-world scenarios, particularly within the Chinese context, whereas the automatically compressed dataset offers a 90% increase in speed and maintains 98% consistency across 18 public benchmarks.
@@
 For automated evaluation, a General Scoring Model is proposed to support 6 kinds of question types with 95% accuracy on OOD models and benchmarks.
-Figure 1: Illustration of the unified ability taxonomy proposed in UNO-Bench.
-(a) exemplarily presents the way to unify uni-modal and omni-modal samples.
-(b) shows the three-layer hierarchical structure of the unified ability taxonomy and its distribution in UNO-Bench.
-Among the 13 abilities, (c) demonstrates the $4$ diverse task types that cover a large range of real world scenarios.
+Dataset Omni-modal Uni-modal Acc. Solvable Source #Tasks #QA Pairs QA Type Language
+MMBench I - - 80% private 20 3217 MC EN/CH
+MMAU A - - 15% private 27 10000 MC EN
+MVBench V - - public 20 4000 MC EN
+OmniBench I+A 75% 90% public 8 1142 MC EN
+AV-Odyssey I+V+A 91% 99% public 26 4555 MC EN
+WorldSense V+A 99% 23% public 26 3172 MC EN
+Daily-Omni V+A 94% 59% public 6 1197 MC EN
+UNO-Bench-omni I+V+A -  100% 98% 90% private 44 1250 MC/MO EN/CH
+UNO-Bench-uni - I/V/A - - 40% private 44 2480 MC EN/CH
+Table 1: Comparison of MultiModal Benchmarks, with I, A, V, and T representing image, audio, video, and text modalities, respectively.
+It reports on the accuracy of question-answer pairs and the percentage of questions requiring omni-modal solutions, labeled as Acc.
+and Solvable.
+The Source category specifies the origin of the materials.
+Private sources, as opposed to public ones, can prevent data contamination.
+QA types include MC for multi-choice questions and MO for multi-step open-ended questions.
+EN and CH denote English and Chinese languages.
+UNO-Bench includes 1250 human-curated samples in the omni-modal section (referred to as -omni) and 2480 enhanced samples in the uni-modal section (referred to as -uni).
 2 Related Work
@@
 Based on large language models, vision language models (VLM) (Bai et al., 2025; Xiaomi, 2025; Zeng et al., 2025) and audio language models (ALM) (Ding et al., 2025; Wu et al., 2025) introduce the general intelligence to vision modality and audio modality respectively.
-Various uni-modal benchmarks conduct comprehensive evaluations on VLMs (Liu et al., 2024b; Lu et al., 2024a; Wang et al., 2024b, c; Liu et al., 2024d; Mathew et al., 2021; Ouyang et al., 2024; Li et al., 2024a; Wu et al., 2024; Liu et al., 2024c; xAI, 2023; Xiao et al., 2021; Huang et al., 2025; Hu et al., 2025; Fu et al., 2024) and ALMs (Ardila et al., 2019; Wang et al., 2021; Yang et al., 2024; Ao et al., 2024).
-For image modality, MMBench(Liu et al., 2024b) proposed a systematically designed benchmark to evaluate general image understanding on 20 different tasks.
-Focused on mathematic, MathVision(Wang et al., 2024b) collected questions from 19 mathematic competitions to evaluate VLMs complex reasoning ability.
-In addition to above, OCRBench(Liu et al., 2024d) supplied the evaluation on text recognition and document understanding.
-For video modality, MVBench(Li et al., 2024a) aggregated 11 public video benchmarks and incorporated data enhancement process to cover 20 dynamic video understanding tasks.
+Various uni-modal benchmarks conduct comprehensive evaluations on VLMs (Liu et al., 2024a; Lu et al., 2024a; Wang et al., 2024a, b; Liu et al., 2024b; Mathew et al., 2021; Ouyang et al., 2024; Li et al., 2024b; Wu et al., 2024; Liu et al., 2024c; xAI, 2023; Xiao et al., 2021; Huang et al., 2025; Hu et al., 2025; Fu et al., 2024) and ALMs (Ardila et al., 2019; Wang et al., 2021; Yang et al., 2024; Ao et al., 2024).
+For image modality, MMBench(Liu et al., 2024a) proposed a systematically designed benchmark to evaluate general image understanding on 20 different tasks.
+Focused on mathematic, MathVision(Wang et al., 2024a) collected questions from 19 mathematic competitions to evaluate VLMs complex reasoning ability.
+In addition to above, OCRBench(Liu et al., 2024b) supplied the evaluation on text recognition and document understanding.
+For video modality, MVBench(Li et al., 2024b) aggregated 11 public video benchmarks and incorporated data enhancement process to cover 20 dynamic video understanding tasks.
 To complement the long video understanding area, LongVideoBench(Wu et al., 2024) introduces hourly video materials to evaluate the information retrieval ability from long context.
 For audio modality, MMAU(Sakshi et al., 2025) provides general audio understanding assessment across speech, sounds and music domains, featuring diverse audio samples.
-There’s massive uni-modal benchmarks covering diverse model abilities on vision modality and audio modality separately.
+There are massive uni-modal benchmarks covering diverse model abilities on vision modality and audio modality separately.
 2.2 Omni-Modal Benchmarks
-Omni models have arisen in recent years(Comanici et al., 2025; Xu et al., 2025b; AI et al., 2025; Li et al., 2025), as the pioneer, Gemini(Comanici et al., 2025) shows a strong ability in understanding both vision and audio, while Qwen-3-Omni(Xu et al., 2025b) provides leading performance in open-source models.
-However, there’s less omni-modal benchmarks that can evaluate the modality combination across image, video and audio.
-OmniBench(Li et al., 2024b) inserted audio as a context into the image understanding task and made up an omni-modal benchmark, while the data quality needs further improvement.
-WorldSence(Hong et al., 2025) emphasized audio-visual data in real world scenarios with high data quality, while most audio-visual questions can be solved by audio or video alone, which can’t assess the cross-modality ability.
+Omni models have arisen in recent years(Comanici et al., 2025; Xu et al., 2025a; AI et al., 2025; Li et al., 2025), as the pioneer, Gemini(Comanici et al., 2025) shows a strong ability in understanding both vision and audio, while Qwen-3-Omni(Xu et al., 2025a) provides leading performance in open-source models.
+However, there are less omni-modal benchmarks that can evaluate the modality combination across image, video and audio.
+OmniBench(Li et al., 2024a) inserted audio as a context into the image understanding task and made up an omni-modal benchmark, while the data quality needs further improvement.
+WorldSence(Hong et al., 2025) emphasized audio-visual data in real world scenarios with high data quality, while most audio-visual questions can be solved by audio or video alone, which cannot assess the cross-modality ability.
 Other datasets focus on audio (Gong et al., 2024) or video (Zhou et al., 2025) and cover limited task types.
+For instance, in Figure.3(b), the problem can be resolved using either the audio modality or the visual modality, whereas in Figure.3(c), only the visual modality is necessary to address the problem.
+These instances are likely to exaggerate the capabilities of the omni model, making it crucial to evaluate the cross-modality solvable problem (illustrated in Figure.3(a)) to accurately assess omni-modal capability (refer to Section.4.3 for more details).
 Addressing these limitations, we propose a novel and unified benchmark, UNO-Bench, that enables comprehensive model assessment and pushing omni model to the next-level of intelligence.
@@
 In this section, we first introduce the omni-modal dataset construction pipeline in Section.3.1.
-The uni-modal dataset improvement method and a general evaluation dataset compression method to improve the evaluation efficiency are introduced in Section.3.2.
-Finally, the multi-step open-ended questions are introduced in Section.3.3.
+For uni-modal dataset, an quality improvement method and a general dataset compression method to improve the evaluation efficiency are introduced in Section.3.2.
+Finally, the novel multi-step open-ended questions are presented alongside a general scoring model in Section.3.3
 3.1 Omni-modal Dataset Construction
 We have established a human-centric data construction pipeline (Figure.4) that efficiently empower human intelligence to produce high-quality and high-diversity dataset.
-Figure 2: Dataset Construction Pipeline includes human crafted process (left side) and automated data compression (right side).
-First, we collect diverse and novel materials to prevent data contamination.
-Second, with the proposed unified ability taxonomy, human annotators including experts will craft questions, answers and record audios in real-world scenarios.
-Finally, with model checking, ablation study and human experts revision, we achieves high quality and diversity dataset.
-Regarding automated data compression, we use model performances as the sample feature, employing clustering and resampling to achieve a 90% dataset compression.
+Figure 2: Illustration of the unified ability taxonomy proposed in UNO-Bench.
 3.1.1 Model Ability Taxonomy
 Through cumulative experiences on multimodal evaluation from both model-side and user-side, we summarize the capabilities of uni-modal and omni-modal into a unified model ability taxonomy.
-As shown in Figure.1, the omni model’s capabilities are systematically categorized into two primary dimensions: Perception and Reasoning.
-Detailed definitions and examples can be found in the Appendix.E.
-Perception dimension structured through seven recognition types: Object Perception, Attribute Perception, Scenario Perception, Spatial Perception, Cross-Modal Conversion, Semantic Understanding.
-In addition, we incorporate Cross-Modal Alignment evaluation to assess information synchronization across modalities.
-The perception dimension fuses visual and audio modalities. For example, Scenario Perception includes the recognition of visual scenes and the judgment of audio scenes.
-Reasoning dimension extends conventional reasoning categories (General, STEM, Code) with: Spatial Reasoning (including Static Reasoning and Dynamic Reasoning), Temporal Reasoning, and Complex Reasoning (multi-conditional, multi-step problem solving).
-The unified ability taxonomy combines uni-modal and omni-modal abilities which provides a comprehensive measurement that is particularly critical for advanced omni models.
-Based on this taxonomy, we create a diversity dataset with 44 task types distributed as shown in Figure.1.
+As shown in Figure.2(b), the omni model’s capabilities are systematically categorized into two primary dimensions: Perception and Reasoning.
+Detailed definitions and examples can be found in the Appendix.D.
+Perception dimension structured through seven recognition types including Object Perception, Attribute Perception, Scenario Perception, Spatial Perception, Cross-Modal Conversion, Semantic Understanding.
+In addition, we incorporate Cross-Modal Alignment to assess information synchronization across modalities.
+Reasoning dimension extends conventional reasoning categories (including General, STEM, Code) with Spatial Reasoning (including Static Reasoning and Dynamic Reasoning), Temporal Reasoning, and Complex Reasoning (which indicates multi-conditional, multi-step problem).
+As shown in Figure.2(a), the unified ability taxonomy combines uni-modal and omni-modal abilities which provides a comprehensive measurement that is particularly critical for omni models.
+For example, Scenario Perception includes the recognition of visual scenes and the judgment of audio scenes.
+Based on this taxonomy, we create a diversity dataset with 44 task types illustrated in Figure.2(c).
 3.1.2 Material Collection
-In both data quality checks and experimental results, we found that natural audio-visual synchronized video data contains a large amount of information redundancy, only a few videos require both audio and vision modality simultaneously.
+In both data quality checks and experimental results, we found that the natural video with audio-visual synchronized data contains a large amount of information redundancy, only a few videos require both audio and visual modality simultaneously.
 Therefore, we begin with carefully designed material collection.
-Our materials have the following three characteristics: a.
+Our materials have the following three characteristics:
 Diverse Sources.
-The majority of our materials are real-world photos and videos collected through crowdsourcing, and another portion sourced from copyright-free website.
+The majority of our materials are real-world photos and videos collected through crowd sourcing, and another portion sourced from copyright-free websites.
 Additionally, a small fraction comes from high-quality public datasets such as MMVU(Zhao et al., 2025), LongVideoBench(Wu et al., 2024), and VideoVista(Chen et al., 2025).
-b.
-Rich and Diverse Topics.
-Our materials cover a broad spectrum of subjects, including society, culture, art, life, literature, science, and so on.
-c.
+Rich and Diverse Topics. Our materials cover a broad spectrum of subjects, including society, culture, art, life, literature, science, and so on.
 Live-Recorded Audio.
 Apart from background sounds and music, all dialogue is recorded by human speakers.
-With over 20 participants in the recording process, the audio features are rich and closely reflect the diverse vocal characteristics of the real world.
+With over 20 participants in the recording process, the audio features are rich and closely reflect the diverse vocal characteristics of the real world, such as Mandarin and Sichuan dialect.
 Finally, we conduct material filtering.
@@
 Additionally, label the materials with more detailed information such as subject, event, scene, and style to facilitate subsequent annotators to quickly find matching materials.
+Figure 3: Illustration of the cross-modality solvable sample.
 3.1.3 QA Annotation
@@
 Human experts have extensive experience in cross-modal data construction and annotation, a deeper understanding of model capabilities, and thus ensure higher professionalism and specificity in the data they construct.
-Most crowd-sourced users are college students with rich experience in multi-modal model interactions and diverse professional backgrounds, providing data with better authenticity and diversity.
+Most crowd-sourced users are college students with rich experience in multimodal model interactions and diverse professional backgrounds, providing data with better authenticity and diversity.
 First, annotators clarify the required image/video features based on task type definitions and filter appropriate materials from existing libraries using tags.
@@
 Through this workflow, we ultimately generate complete QA pairs encompassing three modalities: visual, auditory, and textual.
-Compared to conventional methods limited to human intervention only during the quality assurance phase, our pipeline integrates a "Human-in-the-Loop" approach, ensuring continuous manual involvement from the initial data sourcing to the final output.
+Compared to conventional methods limited to human intervention only during the quality assurance phase, our pipeline integrates a human-centric approach, ensuring continuous manual involvement from the initial data sourcing to the final output.
 This methodology not only prevents data leakage but also more accurately simulates real-world scenarios.
@@
 Each question undergoes at least three rounds of independent quality inspection to maximize data quality.
-First, a preliminary model check is conducted to filter out cases with ambiguous questions, non-unique answers, or those that do not conform to the task type.
-Second, through modality ablation experiments, we remove one modality of information from the QA pair to see if the model can answer based solely on the remaining information.
+Model Check, a preliminary model check is conducted to filter out cases with ambiguous questions, non-unique answers, or those that do not conform to the task type.
+Ablation Study, through modality ablation experiments, we remove one modality of information from the QA pair to see if the model can answer based solely on the remaining information.
 If the question becomes unsolvable or ambiguous after removing any one modality, it proves the cross-modality solvability of the data.
-Finally, manual quality inspection and revision are performed.
+Human Check, finally manual quality inspection and revision are performed.
+Figure 4: Dataset Construction Pipeline includes human-centric process (left side) and automated data compression (right side).
+First, we collect diverse and novel materials to prevent data contamination.
+Second, with the proposed unified ability taxonomy, human annotators including experts will craft questions, answers and record audios in real-world scenarios.
+Finally, with model checking, ablation study and human experts revision, we achieves high quality and diversity dataset.
+Regarding automated data compression, we present a clustering-guided hierarchical sampling method to achieve efficient compression while maintaining high evaluation consistency.
 3.2 Uni-modal Dataset Improvement
@@
 Existing public uni-modal datasets are bothered by data leakage issue(Xu et al., 2024).
-To verify the influence, we adopt privatization improvement on the widely used public dataset MMBench(Liu et al., 2024b).
+To verify the influence, we adopt privatization improvement on the widely used public dataset MMBench(Liu et al., 2024a).
 As shown in Figure.13, the performance of models have better distinguishability after dataset improvement, reflecting the true capability differences between models.
 Therefore, for uni-modal data, we also follow the aforementioned construction process for self-construction datasets.
-Figure 3: Comparison result of privatization improvement. After improvement, the performances among models are more distinguishable.
 In addition to self-constructed data, we also selected some multimodal data from public datasets to supplement in terms of capability items and data types.
@@
 Regarding the existing large-scale uni-modal benchmarks, to reduce the evaluation cost of large-scale models, we designed a clustering-guided hierarchical sampling (CGHS) method as shown in Figure.4.
-This method achieves efficient compression and high evaluation consistency through the following steps:
+CGHS is a general method for dataset compression, which utilizes model performance metrics as features rather than the content of questions to select important samples that impact model performance.
+For training datasets, CGHS can retain both simple and difficult samples in an unsupervised manner or minimize similar rollout samples in a batch for online policy.
+When it comes to test datasets, CGHS is capable of achieving efficient compression while maintaining high evaluation consistency.
+The introduction of CGHS is outlined in the following steps:
 Question Characterization: Represent each question as an x-dimensional vector, where dimensions correspond to scores from different models on that question.
-Cluster-based Stratification: Utilize the Kmeans++(Arthur and Vassilvitskii, 2007) algorithm to categorize questions into k clusters, each representing a "model performance similar" question type (e.g., easy questions, difficult questions, etc.).
+Cluster-based Stratification: Utilize the Kmeans++(Arthur and Vassilvitskii, 2007) algorithm to categorize questions into k clusters, each representing a ”model performance similar” question type (e.g., easy questions, difficult questions, etc.).
 Hierarchical Sampling: Determine the sample size for each stratum based on cluster size proportions, and construct the final evaluation subset through simple random sampling.
-Validity Verification: Use Margin of Error (MoE) and Root Mean Square Error (RMSE) to measure theoretical and practical errors respectively.
-Verify the consistency between model rankings and scores using Spearman Rank Correlation Coefficient (SRCC) and Pearson Linear Correlation Coefficient (PLCC).
+Validity Verification: To verify the compression performance, we define these metrics: Spearman’s Rank Correlation Coefficient (SRCC) for ranking consistency, Pearson’s Linear Correlation Coefficient (PLCC) for linear value consistency, Root Mean Square Error (RMSE) for numerical precision, Margin of Error (MoE) for quantifying estimation uncertainty, and Confidence Interval Coverage (CIC) for statistical reliability.
 To ensure statistical stability, we repeat the above steps by using 5 random splits and performing 10-fold cross-validation.
@@
 (2019) exist, they are hampered by high implementation difficulty, low accuracy, or poor efficiency.
-Moreover, the prevalence of multiple-choice formats in existing benchmarks is unrepresentative of real-world, open-ended user queries and can mask model deficiencies.
+Moreover, the prevalence of multiple-choice formats in existing benchmarks is unrepresentative of real-world, open-ended user queries and may conceal the weaknesses of models.
 To address these issues, we propose an innovative Multi-Step Open-Ended Question (MO) type, designed for granular and realistic assessment.
@@
 Since the task is to compare the target answer and the predicted answer, we use Qwen3-14B(Yang et al., 2025) as backbone and curate a training dataset as illustrated in Figure.6.
-One of the critical way to improve accuracy is to group questions into finer types and define appropriate criteria for each types.
-See details in Appendix.D.
+One of the critical way to improve accuracy is to group questions into finer types and define appropriate criteria for each types, as shown in Figure.7.
 Through the human-in-the-loop dataset curation, the scoring model achieves 95% accuracy in out-of-distribution models and benchmarks.
@@
 multiple-choice questions), multi-step open-ended questions can effectively observe the ability decay of models in long-chain reasoning, providing a more realistic difficulty for advanced models with stronger discrimination.
-Figure 4: Example of multi-step open-ended questions.
-Figure 5: Training dataset construction and model training pipeline for general scoring model.
-Table 2: Omni-modal benchmark results on multimodal models.
-Model Audio Visual Omni-MC Omni-MO
-Qwen-2.5-Omni-3B 54.40 42.67 27.80 24.76
-MiniCPM-O-2.6 56.50 42.27 28.60 23.76
-Ming-lite-Omni-1.5 58.30 46.28 28.90 25.48
-Baichuan-Omni-1.5 54.10 44.66 29.70 21.04
-Qwen-2.5-Omni-7B 60.20 50.68 32.60 27.72
-Qwen-3-Omni-30B-A3B 79.40 63.29 42.10 37.08
-Gemini-2.0-Flash 70.70 62.76 44.90 38.56
-Gemini-2.5-Flash 79.50 69.54 54.30 47.08
-Gemini-2.5-Pro 88.40 78.67 70.90 57.32
-Table 3: Omni-modal analysis on model ability taxonomy. Cross-modal Recognition refers to the set of other Perception capabilities except Cross-modal Alignment.
-Model Perception Reasoning Overall
-Cross-modal
-Alignment
-Cross-modal
-Recognition
-Overall
-General
-Reasoning
-Temporal
-Reasoning
-Spatial
-Reasoning
-Overall
-Qwen-2.5-Omni-3B 29.84 35.94 33.09 20.65 50.00 20.83 23.98 27.80
-MiniCPM-O-2.6 26.70 30.88 28.92 26.62 42.42 26.67 28.40 28.60
-Ming-lite-Omni-1.5 28.80 35.94 32.60 24.38 43.94 24.17 26.53 28.90
-Baichuan-Omni-1.5 30.89 32.26 31.62 25.87 45.45 28.33 28.57 29.70
-Qwen-2.5-Omni-7B 38.22 36.41 37.25 28.11 43.94 26.67 29.59 32.60
-Qwen-3-Omni-30B-A3B 53.40 45.16 49.02 38.06 53.03 26.67 37.41 42.10
-Gemini-2.0-Flash 43.98 49.77 47.06 45.02 57.58 31.67 43.71 44.90
-Gemini-2.5-Flash 56.02 50.69 53.19 61.44 68.18 27.50 55.27 54.30
-Gemini-2.5-Pro 74.35 70.05 72.06 75.62 84.85 45.00 70.41 70.90
+Figure 5: Construction of Multi-Step Open-Ended Questions.
+Figure 6: Training pipeline for general scoring model.
+Figure 7: Definition of finer question types for general scoring model.
 4 Experiment and Analysis
 4.1 Experiment Setting
-We evaluate omni models that support text, visual, and audio inputs simultaneously, including open-source models: Qwen-3-Omni-30B-A3B-Instruct(Xu et al., 2025b), Qwen-2.5-Omni-3B, Qwen-2.5-Omni-7B(Xu et al., 2025a), Baichuan-Omni-1.5(Li et al., 2025), MiniCPM-O-2.6(Yao et al., 2024), and Ming-lite-Omni-1.5(AI et al., 2025), as well as closed-source models: Gemini-2.5-Pro, Gemini-2.5-Flash, and Gemini-2.0-Flash(Comanici et al., 2025).
-To have a fair comparison between instruct model and thinking model, we adopt similar way in Qwen-3(Xu et al., 2025b) that limits thinking budget to 128 tokens.
+We evaluate omni models that support text, visual, and audio inputs simultaneously, including open-source models: Qwen-3-Omni-30B-A3B-Instruct(Xu et al., 2025a), Qwen-2.5-Omni-3B, Qwen-2.5-Omni-7B(Xu et al., 2025b), Baichuan-Omni-1.5(Li et al., 2025), MiniCPM-O-2.6(Yao et al., 2024), and Ming-lite-Omni-1.5(AI et al., 2025), as well as closed-source models: Gemini-2.5-Pro, Gemini-2.5-Flash, and Gemini-2.0-Flash(Comanici et al., 2025).
+To have a fair comparison between instruct model and thinking model, we adopt similar way in Qwen-3(Xu et al., 2025a) that limits thinking budget to 128 tokens.
 We apply this restriction to Gemini-2.5-Pro and disable the thinking mode for both Gemini-2.5-Flash and Gemini-2.0-Flash.
@@
 In video processing, each model receives raw video and performs frame sampling according to its own sampling strategy.
-In the following sections, we conduct comprehensive experiments on UNO-Bench and try to answer the following questions:
+In the subsequent sections, we perform detailed experiments on UNO-Bench and aim to address the following questions:
   1. 1.
-    How about the performance of current omni models and their shortage abilities?
+    How do current omni models perform, and what are their limitations?
   2. 2.
-    What is the relationship of uni-modal ability and omni-modal ability?
+    How are uni-modal and omni-modal capabilities related?
   3. 3.
-    Could the UNO-Bench provide an effective evaluation for the omni model?
+    Is the UNO-Bench capable of effectively evaluating the omni model?
 4.2 Model Performance
@@
 This indicates that advanced reasoning remains a key differentiator and a primary bottleneck for current multimodal models.
+Table 2: General performance of omni models in UNO-Bench for both uni-modal capability and omni-modal capability, where omni-modal benchmark includes multi-choice questions (Omni-MC) and multi-step open-ended questions (Omni-MO).
+Model Audio Visual Omni-MC Omni-MO
+Qwen-2.5-Omni-3B 54.40 42.67 27.80 24.76
+MiniCPM-O-2.6 56.50 42.27 28.60 23.76
+Ming-lite-Omni-1.5 58.30 46.28 28.90 25.48
+Baichuan-Omni-1.5 54.10 44.66 29.70 21.04
+Qwen-2.5-Omni-7B 60.20 50.68 32.60 27.72
+Qwen-3-Omni-30B-A3B 79.40 63.29 42.10 37.08
+Gemini-2.0-Flash 70.70 62.76 44.90 38.56
+Gemini-2.5-Flash 79.50 69.54 54.30 47.08
+Gemini-2.5-Pro 88.40 78.67 70.90 57.32
+Table 3: Analysis of Omni-MC on ability taxonomy. To simplify the analysis, Cross-modal Recognition refers to the set of other Perception capabilities except Cross-modal Alignment.
+Model Perception Reasoning Overall
+Cross-modal
+Alignment
+Cross-modal
+Recognition
+Overall
+General
+Reasoning
+Temporal
+Reasoning
+Spatial
+Reasoning
+Overall
+Qwen-2.5-Omni-3B 29.84 35.94 33.09 20.65 50.00 20.83 23.98 27.80
+MiniCPM-O-2.6 26.70 30.88 28.92 26.62 42.42 26.67 28.40 28.60
+Ming-lite-Omni-1.5 28.80 35.94 32.60 24.38 43.94 24.17 26.53 28.90
+Baichuan-Omni-1.5 30.89 32.26 31.62 25.87 45.45 28.33 28.57 29.70
+Qwen-2.5-Omni-7B 38.22 36.41 37.25 28.11 43.94 26.67 29.59 32.60
+Qwen-3-Omni-30B-A3B 53.40 45.16 49.02 38.06 53.03 26.67 37.41 42.10
+Gemini-2.0-Flash 43.98 49.77 47.06 45.02 57.58 31.67 43.71 44.90
+Gemini-2.5-Flash 56.02 50.69 53.19 61.44 68.18 27.50 55.27 54.30
+Gemini-2.5-Pro 74.35 70.05 72.06 75.62 84.85 45.00 70.41 70.90
 4.2.2 Top-tier Analysis
-Why Gemini-2.5-Pro performances significantly better than other models?
-We attempt to provide an analysis and some hypotheses.
+What makes the performance of Gemini-2.5-Pro stand out compared to other models?
+We aim to offer an analysis along with several hypotheses.
 On one hand, it stems from the leading uni-modal understanding ability.
@@
 Figure 8: An example of omni-modal evaluation result. Gemini-2.5-Pro displays audio captions to assist in solving the problem.
-However, does Gemini performance good enough?
-To answer this question, we invited human experts to conduct a competition.
-It is important to note that, unlike the dataset annotators, the invited human experts have not been exposed to the question or the answer.
+Figure 9: The competition between human experts and Gemini-2.5-Pro. Gemini-2.5-Pro shows comparable perception capability but lower reasoning capability.
+The successive question is whether Gemini’s performance measures up.
+To answer this question, we invited human experts for a competition.
+It’s crucial to highlight that, unlike the dataset annotators, these experts had no prior exposure to the questions or answers.
 Finding 1.
-Gemini-2.5-Pro has reached human comparable perception ability in omni-modal understanding, yet there remains a gap in its reasoning performance.
-Compared to human experts, Gemini-2.5-Pro only has 8.3% less performance, which shows humanity comparable intelligence.
+Gemini-2.5-Pro has reached human comparable perception ability in omni-modal perception, yet there remains a gap in its reasoning performance.
+Compared to human experts, Gemini-2.5-Pro exhibits similar performance in perception, but falls behind in reasoning.
 The comparison of scores for specific ability items can be seen in Figure.9.
@@
 By interviewing various human experts, it becomes evident that humans might miss some information in video or audio formats, and their world knowledge is more limited compared to large language models.
-Figure 7: The competition between human experts and Gemini-2.5-Pro. Gemini-2.5-Pro shows comparable perception capability but lower reasoning capability.
 4.3 Uni-Modal v.s. Cross-Modal
@@
 where $f_{\text{A}},f_{\text{V}}$ represent modality independent path contributions, $f_{\text{I}}$ the interaction, and $b$ a baseline performance constant (e.g. random guess).
-We arrive at the following result through rigorous mathematical derivation, and the detailed derivation process is provided in the Appendix B.
+We arrive at the following result through rigorous mathematical derivation, and the detailed derivation process is provided in the Appendix.B.
 $\mathcal{P}_{\text{Omni}}=C\cdot\mathcal{P}_{\text{A}}^{\alpha}\mathcal{P}_{\text{V}}^{\beta}+b$ (2)
 where $C$ is a scaling constant, and exponents $\alpha,\beta$ model the interaction’s elasticity.
-We then posit a fusion symmetry assumption: in end-to-end omni models, the fusion mechanism does not inherently favor one modality over anotherXu et al.
-(2025b); Yao et al.
-(2024), implying symmetric scaling behavior.
+We then posit a fusion symmetry assumption: in end-to-end omni models, the fusion mechanism does not inherently favor one modality over another (Xu et al., 2025a; Yao et al., 2024), implying symmetric scaling behavior.
 This leads to $\alpha=\beta$.
 Substituting this into Eq.
-10, we arrive at the Omni-modal Compositional Law:
+2, we arrive at the Omni-modal Compositional Law:
 $\mathcal{P}_{\text{Omni}}=C\cdot(\mathcal{P}_{\text{A}}\times\mathcal{P}_{\text{V}})^{\alpha}+b$ (3)
@@
 Because $\alpha>1$, the function is convex, meaning the rate of performance gain accelerates.
-This explains the transition from a "short-board effect" to an "emergent ability" seen in Figure.8:
+This explains the transition from a ”short-board effect” to an ”emergent ability” seen in Figure.10:
   • Limited Gains at Low Performance: For models with weaker uni-modal abilities (e.g., MiniCPM-O), the curve is relatively flat.
 Small improvements in the product of uni-modal scores yield only marginal gains in omni-modal performance.
-This can be seen as a "short-board effect", where the system is not yet capable of effectively leveraging the combined inputs.
+This can be seen as a ”short-board effect”, where the system is not yet capable of effectively leveraging the combined inputs.
   • Emergent Ability at High Performance: As uni-modal abilities strengthen (e.g., moving towards Gemini-2.5-Pro), the curve steepens dramatically.
@@
 Next, we conduct ablation studies to dig dive about the enhancement from vision and audio modality respectively.
-Table 4: Ablation of visual understanding ability.
-Model Perception Reasoning Overall
-Audio + Caption + Visual Audio + Caption + Visual Audio + Caption + Visual
-Qwen-2.5-Omni-3B 17.76 29.13 33.09 20.07 21.43 23.98 19.12 24.60 27.80
-MiniCPM-O-2.6 29.44 29.61 28.92 27.21 29.93 28.40 28.13 29.80 28.60
-Ming-lite-Omni-1.5 26.28 31.07 32.60 23.13 21.43 26.53 24.42 25.40 28.90
-Baichuan-Omni-1.5 22.14 32.04 31.62 23.81 26.70 28.57 23.12 28.90 29.70
-Qwen-2.5-Omni-7B 22.14 30.10 37.25 20.41 25.34 29.59 21.12 27.30 32.60
-Qwen-3-Omni-30B-A3B 27.01 46.84 49.02 18.71 39.63 37.41 22.12 42.60 42.10
-Gemini-2.0-Flash 25.55 44.17 47.06 29.76 45.58 43.71 28.03 45.00 44.90
-Gemini-2.5-Flash 22.63 49.03 53.19 29.08 53.23 55.27 26.43 51.50 54.30
-Gemini-2.5-Pro 37.71 63.83 72.06 42.18 65.99 70.41 40.34 65.10 70.90
 4.3.1 Ablation Visual Understanding
@@
 This suggests that for these models, the language processing pathway may be more adept at extracting semantic meaning than their own visual encoders, highlighting a potential imbalance in their multimodal processing capabilities.
+Table 4: Ablation of visual understanding ability.
+The three settings are audio-only (Audio), audio plus high-quality textual captions of the visual scene (+Caption), and the full audio-visual input (+Visual).
+Model Perception Reasoning Overall
+Audio + Caption + Visual Audio + Caption + Visual Audio + Caption + Visual
+Qwen-2.5-Omni-3B 17.76 29.13 33.09 20.07 21.43 23.98 19.12 24.60 27.80
+MiniCPM-O-2.6 29.44 29.61 28.92 27.21 29.93 28.40 28.13 29.80 28.60
+Ming-lite-Omni-1.5 26.28 31.07 32.60 23.13 21.43 26.53 24.42 25.40 28.90
+Baichuan-Omni-1.5 22.14 32.04 31.62 23.81 26.70 28.57 23.12 28.90 29.70
+Qwen-2.5-Omni-7B 22.14 30.10 37.25 20.41 25.34 29.59 21.12 27.30 32.60
+Qwen-3-Omni-30B-A3B 27.01 46.84 49.02 18.71 39.63 37.41 22.12 42.60 42.10
+Gemini-2.0-Flash 25.55 44.17 47.06 29.76 45.58 43.71 28.03 45.00 44.90
+Gemini-2.5-Flash 22.63 49.03 53.19 29.08 53.23 55.27 26.43 51.50 54.30
+Gemini-2.5-Pro 37.71 63.83 72.06 42.18 65.99 70.41 40.34 65.10 70.90
 4.3.2 Ablation Audio Understanding
 To isolate the impact of auditory information, we evaluated models under three conditions: visual-only (Visual), visual plus transcribed audio (+Caption), and the full audio-visual input (+Audio).
-We further divided the audio into two categories: Environment sounds, for which we provided textual descriptions, and Speech, for which we provided ASR transcripts.
-All transcriptions were generated by a powerful multimodal model.
+We further divided the audio into three categories: the Speech category was annotated with ASR transcripts, while both the Environment and Music categories received textual descriptions.
+To ensure the robustness of our analysis and improve statistical reliability, the data-insufficient Music class was merged with the Environment class.
+The majority of the transcriptions were manually produced by human annotators, while a smaller subset was generated by a powerful multimodal model.
 The results are presented in Table.5.
@@
 Table 5: Ablation of audio understanding ability.
+The three settings are visual-only (Visual), visual plus transcribed audio (+Caption), and the full audio-visual input (+Audio).
+We further divided the audio into two categories: Environment sounds, for which we provided textual descriptions, and Speech, for which we provided ASR transcripts.
 Model Environment Speech Overall
@@
 To eliminate the random factor, we conduct 5-fold settings and evaluate 10 times on each setting.
-To verify the compression performance, we define these metrics: Spearman’s Rank Correlation Coefficient(SRCC) for ranking consistency, Pearson’s Linear Correlation Coefficient(PLCC) for linear value consistency, Root Mean Square Error(RMSE) for numerical precision, Margin of Error(MoE) for quantifying estimation uncertainty, and Confidence Interval Coverage(CIC) for statistical reliability.
 The experimental result is shown in Figure.11.
@@
 As shown in Table.1, UNO-Bench has 100% accuracy on omni-modal dataset while 98% questions requires cross-modality to solve.
-It shows the highest quality among existing omni model benchmarks.
+It shows the highest quality among existing omni benchmarks.
 An effective benchmark must provide both a clear performance ladder and a meaningful difficulty range, and our UNO-Bench is engineered to deliver on both fronts as shown in Figure.13.
@@
 By combining a structured performance ladder with a balanced challenge, UNO-Bench serves as a more reliable and insightful tool for gauging genuine progress in the field.
+Figure 12: Comparison result of privatization improvement. After improvement, the performances among models are more distinguishable.
 Figure 13: Comparison between omni benchmarks.
 5 Conclusion
-In this work, we introduce a high quality and diversity benchmark to evaluate omni models comprehensively. The main foundings are summarized as below.
-1.
-With unified data framework in UNO-Bench, we found that the multimodal capability may not simply be a linear superposition of uni-modalities, but rather follows a significant multiplicative relationship.
+In this work, we introduce a high quality and diversity benchmark to evaluate omni models comprehensively.
+With unified data framework in UNO-Bench, we found that the omni-modal capability may not simply be a linear superposition of uni-modal capabilities, but rather follows a significant multiplicative relationship.
 The evaluation results show that it manifests as a bottleneck effect on weak models, while exhibiting synergistic promotion on strong models.
-2.
-Both uni-modal and omni-modal understanding capability of the Gemini series far surpasses existing open-source omni models.
+In addition, we found that both uni-modal and omni-modal understanding capability of the Gemini series far surpasses existing open-source omni models.
 The Gemini-2.5-Pro shows comparable perception capability with human experts but still has a performance gap in reasoning aspect.
-3. Besides better dataset quality and evaluation efficiency, UNO-Bench can provide sufficient metric discriminability and a progressive difficulty scale to drive model capability growth.
-In the future work, we will extend the dataset’s scale by human-in-the-loop automated pipeline and hold a private test set to avoid hacking.
+Besides better dataset quality and evaluation efficiency, UNO-Bench can provide sufficient metric discriminability and a progressive difficulty scale to drive model capability growth.
+In the future work, we will extend the dataset’s scale by the human-in-the-loop automated pipeline and hold a private test set to avoid hacking.
 The ability coverage also needs to extend to more difficult reasoning tasks like STEM and code.
 At the same time, the relationship among cross-modals understandings and how to improve them are still exciting problem to explore.
-Furthermore, our compositional law has only been validated on UNO-Bench with uniformly distributed capabilities.
+Furthermore, our compositional law has been validated on UNO-Bench with uniformly distributed capabilities.
 Whether this law still holds under different capability distributions remains to be explored.
@@
 In addition to publicly available benchmarks, we incorporated several in-house benchmarks both before and after compression to ensure the reasonableness of the data distribution.
-  • General visual question answering, RealWorldQAxAI (2023), MMEChaoyou et al. (2023), SeedBenchLi et al. (2023).
-  • Document and chart understanding, OCRBench Liu et al. (2024d), FoxLiu et al. (2024a), DocLocal4kYe et al. (2023).
+  • General visual question answering, RealWorldQAxAI [2023], MMEChaoyou et al. [2023], SeedBenchLi et al. [2023].
+  • Document and chart understanding, OCRBench Liu et al. [2024b], FoxLiu et al. [2024d], DocLocal4kYe et al. [2023].
   • Stem & reasoning, MMMUYue et al.
-(2024), MMMU-ProYue et al.
-(2025), CMMMUGe et al.
-(2024), MathVistaLu et al.
-(2024b), MathVisionWang et al.
-(2024b), ScienceVistaTeam et al.
-(2025), GMAI-MMBenchYe et al.
-(2024).
-  • Multi-image Understanding, ReMiKazemi et al. (2024), MuirBenchWang et al. (2024a).
-  • Audio question answering, MMAUSakshi et al. (2025), MMSUWang et al. (2025), SDQAFaisal et al. (2021).
+[2024], MMMU-ProYue et al.
+[2025], CMMMUGe et al.
+[2024], MathVistaLu et al.
+[2024b], MathVisionWang et al.
+[2024a], ScienceVistaTeam et al.
+[2025], GMAI-MMBenchYe et al.
+[2024].
+  • Multi-image Understanding, ReMiKazemi et al. [2024], MuirBenchWang et al. [2024c].
+  • Audio question answering, MMAUSakshi et al. [2025], MMSUWang et al. [2025], SDQAFaisal et al. [2021].
 Appendix B Rigorous Derivation of the Compositional Law
@@
 This hints that while our symmetric law captures the primary collaborative effect, the visual component may play a slightly more dominant role, a direction for future investigation.
-Appendix D General Scoring Model
-One of the critical way to improve accuracy is to group questions into finer types and define appropriate criteria for each types. We define six different question types in Figure.12.
-Figure 12: Definition of finer question types for general scoring model.
-Appendix E Model Ability Taxonomy
+Appendix D Model Ability Taxonomy
 This section will provide specific definitions for each ability item and present examples of various task types.
@@
 Figure 18: Example of each ability in reason dimension.
-Appendix F Omni-modal Solubility and Uni-modal Solubility Cases
-Omni-modal solubility means that both visual and audio information are required to answer the question.
-While uni-modal solubility means that only visual or audio information is required to answer the question.
-There are two uni-modal solubility cases.
-One is when audio and video present the same amount of information, and only one modality is needed to solve the problem, as shown in Figure.17(b).
-The other is when one modality provides useful information, while the other modality is unhelpful, as shown in Figure.17(c).
-The omni-modal solubility case can be seen in Figure.17(a).
-Figure 17: Comparison of the problem solvable by each modality.